### v1

# Spectral gap bounds for reversible hybrid Gibbs chains

Abstract
Hybrid Gibbs samplers represent a prominent class of approximated Gibbs algorithms that utilize Markov chains to approximate conditional distributions, with the Metropolis-within-Gibbs algorithm standing out as a well-known example. Despite their widespread use in both statistical and non-statistical applications, very little is known about their convergence properties. This article introduces novel methods for establishing bounds on the convergence rates of hybrid Gibbs samplers. In particular, we examine the convergence characteristics of hybrid random-scan Gibbs and data augmentation algorithms. Our analysis reveals that the absolute spectral gap of a reversible hybrid chain can be bounded based on the absolute spectral gap of the exact Gibbs chain and the absolute spectral gaps of the Markov chains employed for conditional distribution approximations. The new techniques are applied to four algorithms: a random-scan Metropolis-within-Gibbs sampler, a hybrid proximal sampler, random-scan Gibbs samplers with block updates, and a hybrid slice sampler.
Keywords: convergence rate, Metropolis-within-Gibbs, MCMC, proximal sampler, slice sampler.
1 Introduction
The Gibbs sampler (Geman and Geman, 1984; Casella and George, 1992) is an immensely popular Markov chain Monte Carlo (MCMC) algorithm. It can be used to sample from the joint distribution of multiple variables by iteratively updating each variable based on its conditional distribution while keeping the other variables fixed. In practice, the conditional distributions themselves may be difficult to sample from. Algorithms like the Metropolis-within-Gibbs algorithm (Metropolis et al., 1953; Chib and Greenberg, 1995; Tierney, 1998) circumvent this problem by approximating an intractable conditional distribution with a Markovian step that leaves the conditional distribution invariant. We refer to these algorithms as “hybrid” Gibbs algorithms, as opposed to “exact” Gibbs algorithms that make perfect draws from conditional distributions. Since their inception, hybrid Gibbs samplers have been exploited to solve many complex problems from various domains, including pharmacokinetics (Gilks et al., 1995), biomedical informatics (Wang et al., 2023), econometrics (Schamberger et al., 2017), deep latent variable models (Mattei and Frellsen, 2018), and privacy-protected data analysis (Ju et al., 2022).
Although these hybrid algorithms have enjoyed pervasive popularity and extensive application, there remains a notable scarcity of analytical techniques tailored for understanding their convergence properties. This article aims to address this gap. Intuitively, if the Markovian steps in a hybrid Gibbs sampler approximate the intractable conditional distributions well, then the convergence rate of the hybrid Gibbs sampler should be similar to that of the corresponding exact Gibbs sampler. Our analysis confirms this intuition and makes this notion precise with some quantitative convergence rate inequalities. In particular, we provide bounds on the absolute spectral gap of the hybrid algorithm in terms of that of the exact algorithm and the quality of the approximation. Bounds involving Dirichlet forms are also provided. We focus on the case where both the exact Gibbs and the hybrid algorithms are reversible. To be precise, we study random-scan Gibbs and data augmentation algorithms. Moreover, the Markovian approximation of a conditional distribution is assumed to be reversible. The efficiency of a reversible MCMC sampler is closely related to its spectral gap. For a Markov chain associated with a Markov transition kernel (Mtk) $K$ that is reversible with respect to a distribution $\omega$, its absolute spectral gap is defined to be $1-\|K\|_{\omega}$, where $\|K\|_{\omega}\in[0,1]$ is the $L^{2}$ norm of $K$ when $K$ is viewed as a Markov operator. (See Section 3 for the exact definitions.) The $L^{2}$ distance between the distribution of the chain’s $t$th iteration and its target $\omega$ decreases at a rate of $\|K\|_{\omega}^{t}$ as $t$ tends to infinity. Indeed, the norm $\|K\|_{\omega}$ can be regarded as the convergence rate of the chain, with a smaller rate indicating faster convergence. The number of iterations it takes for the chain to reach within some $L^{2}$ distance of its stationary distribution is roughly proportional to $-1/\log\|K\|_{\omega}$, which is approximately $1/(1-\|K\|_{\omega})$ when $\|K\|_{\omega}$ is not too close to 0. Our results give quantitative upper and lower bounds on $(1-\|K_{2}\|_{\omega})/(1-\|K_{1}\|_{\omega})$ when $K_{1}$ and $K_{2}$ are the Mtks of a Gibbs sampler and its hybrid version. In Section 5, we apply the spectral gap bounds, derived under general conditions, to four examples, illustrating their relevance and potential for various applications. Together, these examples effectively showcase how our framework enriches and diversifies the analysis of exact and hybrid Gibbs samplers across different scenarios.
  • Firstly, our method can be efficiently integrated with existing techniques to establish new and explicit convergence bounds. In Section 5.1, we study a random-scan Gibbs sampler whose conditional distributions are approximated using a random-walk Metropolis algorithm. We derive an explicit relationship between the spectral gaps of the exact and hybrid samplers when the conditional distributions are well-conditioned using recent studies on log-concave sampling (Andrieu et al., 2022b). In Section 5.2, we study a new type of proximal sampler (Lee et al., 2021a), which can be regarded as a data augmentation algorithm, or a deterministic-scan Gibbs algorithm with two components. We quantitatively bound its convergence rate when one of the conditional distributions involved is approximated via a random-walk Metropolis step.
  • Secondly, our theory can be used to generalize existing frameworks for analyzing Gibbs algorithms. In Section 5.3, we apply our method to analyze random-scan Gibbs samplers that update multiple coordinates at a time. In particular, we provide a quantitative relationship between two samplers that update, respectively, $\ell$ and $\ell^{\prime}$ coordinates at a time, in terms of their spectral gaps. This extends the main result (Theorem 2) in Qin and Wang (2023), which itself is a generalization of a breakthrough technique called “spectral independence” recently developed in theoretical computer science for analyzing the convergence properties of Gibbs algorithms (Anari et al., 2021; Chen et al., 2021; Blanca et al., 2022; Feng et al., 2022).
  • Lastly, our technique may improve existing studies of some well-known problems. In Section 5.4, we revisit a hybrid slice sampler previously studied by Łatuszyński and Rudolf (2014). By viewing it as a hybrid Gibbs sampler with two components, we derive a convergence rate bound that is similar but slightly better than the one obtained in Łatuszyński and Rudolf (2014).
We now review some relevant existing works. On finite state spaces, Liu (1996) constructed a Metropolized random-scan Gibbs sampler and showed that it is asymptotically more efficient than the corresponding exact Gibbs sampler using Peskun’s ordering. Analysis on general state spaces is more challenging. The convergence properties of hybrid Gibbs samplers were studied qualitatively by Roberts and Rosenthal (1997), Fort et al. (2003), and Qin and Jones (2022). There have been very few quantitative studies on the subject. When the approximating Markov chains are uniformly ergodic with convergence rates bounded away from one, some quantitative convergence bounds are given in Roberts and Rosenthal (1998), Neath and Jones (2009), and Jones et al. (2014). The bounds herein rely on the weaker assumption that the approximating chains are geometrically ergodic with convergence rates bounded away from unity. In fact, one of our bounds for hybrid data augmentation chains relies on yet a weaker condition. Some of the linear algebraic techniques that we employ are adopted from Qin and Wang (2023) and Łatuszyński and Rudolf (2014). The rest of this article is organized as follows. An overview of our results is given in Section 2. Section 3 gives a brief survey on the $L^{2}$ theory of Markov chains, and contains the definitions of concepts like Markov operator and spectral gap. The main results are formally stated in Section 4. In Section 5 we apply our main result to four examples. Section 6 contains a brief discussion on avenues for future research. The Appendix contains some technical results and proofs.
2 Overview
In this section, we describe two hybrid Gibbs samplers of interest, and give an overview of our main results. The first hybrid sampler simulates a random-scan Gibbs chain that involves $n$ possibly approximated conditional distributions. The second is a data augmentation algorithm that simulates a marginal deterministic-scan Gibbs chain with two components. In the second sampler, one of the two conditional distributions involved is approximated.
2.1 Gibbs and hybrid Gibbs samplers
We begin by formally defining conditional distributions through the notion of disintegration. Let $(\mathsf{X},\mathcal{B},\Pi)$ be a probability space. For $i=1,\dots,n$ where $n$ is a positive integer, let $(\mathsf{Y}_{i},\mathcal{F}_{i})$ be a measurable space, and let $\psi_{i}:\mathsf{X}\to\mathsf{Y}_{i}$ be a measurable function. Let $M_{i}=\Pi\circ\psi_{i}^{-1}$, and let $(P_{i,y_{i}})_{y_{i}\in\mathsf{Y}_{i}}$ be a disintegration of $\Pi$ with respect to $\psi_{i}$. That is, (i) for $y_{i}\in\mathsf{Y}_{i}$, $P_{i,y_{i}}(\cdot)$ is a $\sigma$-finite measure, and, for $M_{i}$-a.e. $y_{i}\in\mathsf{Y}_{i}$, $P_{i,y_{i}}(\cdot)$ is a probability measure concentrated on the pre-image $\psi^{-1}_{i}(y_{i}):=\psi^{-1}_{i}(\{y_{i}\})=\{x\in\mathsf{X}:\,\psi_{i}(x)=y_{i}\}$, (ii) for $f:\mathsf{X}\to[0,\infty)$ measurable, $y_{i}\mapsto\int_{\mathsf{X}}f(x)\,P_{i,y_{i}}(\mathrm{d}x)$ is measurable, and
$\int_{\mathsf{X}}f(x)\,\Pi(\mathrm{d}x)=\int_{\mathsf{Y}_{i}}\int_{\mathsf{X}}f(x)\,P_{i,y_{i}}(\mathrm{d}x)\,M_{i}(\mathrm{d}y_{i}).$ (1)
If $X\sim\Pi$, then $M_{i}$ is the marginal distribution of $\psi_{i}(X)$, and $P_{i,y_{i}}(\cdot)$ is a version of the conditional distribution of $X$ given $\psi_{i}(X)=y_{i}$. If $(\mathsf{X},\mathcal{B})$ is a Polish measurable space (i.e., $\mathsf{X}$ is a separable completely metrizable topological space and $\mathcal{B}$ is its Borel algebra), while $\mathcal{B}_{i}$ is countably generated and contains all singletons, then such a disintegration exists and is unique up to an $M_{i}$ almost sure equivalence (see, e.g., Chang and Pollard, 1997).
Example 1 (Joint and conditional distributions on a poduct space).
Suppose that $(\mathsf{X},\mathcal{B})$ is the product of $n$ measurable spaces, say $((\mathsf{X}_{i},\mathcal{B}_{i}))_{i=1}^{n}$. Let $(\mathsf{Y}_{i},\mathcal{F}_{i})$ be the product of the spaces from $(\mathsf{X}_{1},\mathcal{B}_{1})$ to $(\mathsf{X}_{n},\mathcal{B}_{n})$ excluding $(\mathsf{X}_{i},\mathcal{B}_{i})$, and let $\psi_{i}((x_{1},\dots,x_{n}))=x_{-i}$, where $x_{-i}=(x_{1},\dots,x_{i-1},x_{i+1},\dots,x_{n})$. Then, for $X=(X_{1},\dots,X_{n})\sim\Pi$, $M_{i}$ gives the marginal distribution of $X_{-i}:=\psi_{i}(X)$, and $P_{i,y_{i}}(\cdot)$ gives the conditional distribution of $X$ given $X_{-i}=y_{i}$. In practice, it is perhaps more common to consider the conditional distribution of $X_{i}$ given $X_{-i}=y_{i}$, which we shall denote by $\Pi_{i,y_{i}}(\cdot)$. To be formal, we assume that $\Pi_{i,y_{i}}(\cdot)$ is a measure on $(\mathsf{X}_{i},\mathcal{B}_{i})$ that forms a disintegration of the form $P_{i,y_{i}}(\mathrm{d}(x_{1},\dots,x_{n}))=\Pi_{i,y_{i}}(\mathrm{d}x_{i})\,\delta_{y_{i}}(\mathrm{d}x_{-i})$, where $\delta_{y_{i}}$ means the point mass at $y_{i}$. Drawing $x=(x_{1},\dots,x_{n})$ from $P_{i,y_{i}}(\cdot)$ is tantamount to drawing $x_{i}$ from $\Pi_{i,y_{i}}(\cdot)$ and letting $x_{-i}=y_{i}$.
The following lemma, which can be proved using (1) and the concentration of $P_{i,y_{i}}(\cdot)$ on $\psi^{-1}_{i}(y_{i})$, will be useful.
Lemma 2.
For $i=1,\dots,n$, the following holds:
  1. (i)
    For $\Pi$-a.e. $x\in\mathsf{X}$ and $A\in\mathcal{B}$,
    $\int_{\mathsf{X}}P_{i,\psi_{i}(x)}(\mathrm{d}x^{\prime})\,P_{i,\psi_{i}(x^{\prime})}(A)=P_{i,\psi_{i}(x)}(A);$
  2. (ii)
    For $A,B\in\mathcal{B}$,
    $\int_{A}\Pi(\mathrm{d}x)\,P_{i,\psi_{i}(x)}(B)=\int_{B}\Pi(\mathrm{d}x)\,P_{i,\psi_{i}(x)}(A).$
A random-scan Gibbs chain targeting $\Pi$ is a Markov chain with transition kernel $T:\mathsf{X}\times\mathcal{B}\to[0,1]$ given by
$T(x,A)=\sum_{i=1}^{n}p_{i}P_{i,\psi_{i}(x)}(A),$
[…]
  • (H)
    For $i=1,\dots,n$, and $A\in\mathcal{B}$, the set $\{(y_{i},x)\in\mathsf{Y}_{i}\times\mathsf{X}:\,y_{i}=\psi_{i}(x)\}$ is measurable, and the function $(y_{i},x)\mapsto Q_{i,y_{i}}(x,A)$ is measurable.
One can then define a hybrid random-scan Gibbs chain with the Mtk
$\hat{T}(x,A)=\sum_{i=1}^{n}p_{i}Q_{i,\psi_{i}(x)}(x,A).$
In this algorithm, given the current state $x$ and having drawn the index $i$, instead of drawing the next state from the possibly intractable $P_{i,\psi_{i}(x)}(\cdot)$, one simulates one step of the Markov chain associated with the Mtk $(x,A)\mapsto Q_{i,\psi_{i}(x)}(x,A)$ to update the current state.
Example 3 (Markovian approximations on a product space).
Consider again Example 1. Suppose that, in general, $\Pi_{i,y_{i}}(\cdot)$ cannot be efficiently sampled from. However, it is often possible to construct, for each value of $i$ and $y_{i}$, an Mtk $R_{i,y_{i}}:\mathsf{X}_{i}\times\mathcal{B}_{i}\to[0,1]$ that is reversible respect to $\Pi_{i,y_{i}}(\cdot)$. In this case, for $x=(x_{1},\dots,x_{n})$, one can take $Q_{i,y_{i}}(x,\mathrm{d}(x_{1}^{\prime},\dots,x_{n}^{\prime}))=R_{i,y_{i}}(x_{i},\mathrm{d}x_{i}^{\prime})\,\delta_{y_{i}}(\mathrm{d}x^{\prime}_{-i})$. Drawing $(x_{1}^{\prime},\dots,x_{n}^{\prime})$ from $Q_{i,\psi_{i}(x)}(x,\cdot)$ is tantamount to drawing $x_{i}^{\prime}$ from $R_{i,x_{-i}}(x_{i},\cdot)$ and setting $x^{\prime}_{-i}=x_{-i}$. We argue in Section 4.3 that the chains associated with $Q_{i,y_{i}}$ and $R_{i,y_{i}}$ share the same convergence rate.
The following result is standard, and can be proved using (1), the concentration of $P_{i,y_{i}}$ on $\psi^{-1}_{i}(y_{i})$, and the reversibility of $Q_{i,y_{i}}$.
Lemma 4.
For $i=1,\dots,n$, the Mtk $(x,A)\mapsto Q_{i,\psi_{i}(x)}(x,A)$ is reversible with respect to $\Pi$. As a consequence, $\hat{T}$ is reversible with respect to $\Pi$.
Another type of Gibbs chain that we shall study is a marginal deterministic-scan Gibbs chain with two components, which defines a data augmentation algorithm (Tanner and Wong, 1987; van Dyk and Meng, 2001; Hobert, 2011). To be specific, suppose that $n=2$, and let $S:\mathsf{Y}_{1}\times\mathcal{F}_{1}\to[0,1]$ be an Mtk of the form
$S(y_{1},A)=\int_{\mathsf{X}}P_{1,y_{1}}(\mathrm{d}x)P_{2,\psi_{2}(x)}(\psi_{1}^{-1}(A)).$
To simulate this chain, given the current state $y_{1}$, one first draws $x$ from the conditional distribution $P_{1,y_{1}}(\cdot)$, then draws $x^{\prime}$ from the other conditional distribution $P_{2,\psi_{2}(x)}(\cdot)$ and sets the next state to $\psi_{1}(x^{\prime})$. Using Lemma 2 one can show that $S$ is reversible with respect to $M_{1}$.
Example 5 (Standard data augmentation).
Consider Example 1 with $n=2$. One can see that $(\mathsf{Y}_{1},\mathcal{F}_{1})=(\mathsf{X}_{2},\mathcal{B}_{2})$, $(\mathsf{Y}_{2},\mathcal{F}_{2})=(\mathsf{X}_{1},\mathcal{B}_{1})$. Let $X=(X_{1},X_{2})\sim\Pi$. Then $\psi_{1}(X)=X_{-1}=X_{2}$, $\psi_{2}(X)=X_{-2}=X_{1}$, $M_{1}$ is the marginal distribution of $X_{-1}$, and $M_{2}$ is that of $X_{-2}$. The conditional distribution $P_{1,y_{1}}(\mathrm{d}(x_{-2},x_{-1}))=\Pi_{1,y_{1}}(\mathrm{d}x_{-2})\,\delta_{y_{1}}(\mathrm{d}x_{-1})$ gives the conditional distribution of $X$ given $X_{-1}=y_{1}$, while $P_{2,y_{2}}(\mathrm{d}(x_{-2},x_{-1}))=\Pi_{2,y_{2}}(\mathrm{d}x_{-1})\,\delta_{y_{2}}(\mathrm{d}x_{-2})$ gives that of $X$ given $X_{-2}=y_{2}$. The chain associated with $S$ can be simulated as follows: given the current state $y_{1}$, first draw $y_{2}$ using the conditional distribution of $X_{-2}$ given $X_{-1}=y_{1}$, and then draw the next state using the conditional distribution of $X_{-1}$ given $X_{-2}=y_{2}$.
Suppose that one can sample from $P_{1,y_{1}}(\cdot)$ for $y_{1}\in\mathsf{Y}_{1}$, but not from $P_{2,y_{2}}(\cdot)$ for at least some $y_{2}\in\mathsf{Y}_{2}$. However, assume that, for $M_{2}$-a.e. $y_{2}\in\mathsf{Y}_{2}$ and $x\in\psi^{-1}_{2}(y_{2})$, one can sample from $Q_{2,y_{2}}(x,\cdot)$, where $Q_{2,y_{2}}$ is an Mtk that is reversible with respect to $P_{2,y_{2}}(\cdot)$. Then one can simulate the hybrid data augmentation chain with Mtk
$\hat{S}(y_{1},A)=\int_{\mathsf{X}}P_{1,y_{1}}(\mathrm{d}x)\,Q_{2,\psi_{2}(x)}(x,\psi_{1}^{-1}(A)).$
In this algorithm, given the current state $y_{1}$, one first draws $x$ from the conditional distribution $P_{1,y_{1}}(\cdot)$, then updates $x$ to $x^{\prime}$ by simulating one step of the chain associated with the Mtk $(x,A)\mapsto Q_{2,\psi_{2}(x)}(x,A)$, and sets the next state to $\psi_{1}(x^{\prime})$.
Example 6 (Standard data augmentation with an intractable conditional).
Consider again Example 5. Assume that $\Pi_{2,y_{2}}(\cdot)$ cannot be efficiently sampled from. However, it is possible to construct, for each value of $y_{2}$, an Mtk $R_{2,y_{2}}:\mathsf{Y}_{1}\times\mathcal{F}_{1}\to[0,1]$ that is reversible with respect to $\Pi_{2,y_{2}}(\cdot)$. Then, for $x=(x_{-2},x_{-1})$, one can take $Q_{2,y_{2}}(x,\mathrm{d}(x_{-2}^{\prime},x_{-1}^{\prime}))=R_{2,y_{2}}(x_{-1},\mathrm{d}x^{\prime}_{-1})\,\delta_{y_{2}}(\mathrm{d}x^{\prime}_{-2})$. Drawing $(x_{-2}^{\prime},x_{-1}^{\prime})$ from $Q_{2,\psi_{2}(x)}(x,\cdot)$ is tantamount to drawing $x^{\prime}_{-1}$ from $R_{2,x_{-2}}(x_{-1},\cdot)$ and setting $x^{\prime}_{-2}=x_{-2}$. In particular, drawing from $Q_{2,\psi_{2}(x)}(x,\psi_{1}^{-1}(\cdot))$ is equivalent to drawing from $R_{2,x_{-2}}(x_{-1},\cdot)$, so
$\hat{S}(y_{1},A)=\int_{\mathsf{Y}_{2}}\Pi_{1,y_{1}}(\mathrm{d}y_{2})\,R_{2,y_{2}}(y_{1},A).$
Lemma 7.
$\hat{S}$ is reversible with respect to $M_{1}$.
Proof.
For $A,B\in\mathcal{F}_{1}$,
$\int_{A}M_{1}(\mathrm{d}y_{1})\hat{S}(y_{1},B)$
[…]
$=$ $\int_{\mathsf{X}}\Pi(\mathrm{d}x)\,Q_{2,\psi_{2}(x)}(x,\psi_{1}^{-1}(B))\,1_{A}(\psi_{1}(x))\quad(\text{concentration of }P_{1,\psi_{1}(x)}(\cdot))$
[…]
By Lemma 4, the last line is invariant when $A$ and $B$ are exchanged. This establishes the reversibility of $\hat{S}$
2.2 A loose summary of main results
We now give an overview of our main results. The exact statements are given in Section 4, after a brief survey on the $L^{2}$ theory of Markov chains in Section 3. For now, recall that the $L^{2}$ norm of a Markov operator is in $[0,1]$. When the associated chain is reversible, the norm can be viewed as its convergence rate, with a smaller rate indicating faster convergence to the stationary distribution. It is also common to quantify the convergence speed of chains through the spectral gap, which is one minus the operator norm. The norms of $T$, $\hat{T}$, $S$, $\hat{S}$, and $Q_{i,y_{i}}$ are denoted by $\|T\|_{\Pi}$, $\|\hat{T}\|_{\Pi}$, $\|S\|_{\Pi}$, $\|\hat{S}\|_{\Pi}$, and $\|Q_{i,y_{i}}\|_{P_{i,y_{i}}}$, respectively. For the random-scan Gibbs sampler, our bound loosely goes like this:
$\left(1-\sup_{i,y_{i}}\|Q_{i,y_{i}}\|_{P_{i,y_{i}}}\right)(1-\|T\|_{\Pi})\leq 1-\|\hat{T}\|_{\Pi}\leq\left(1+\sup_{i,y_{i}}\|Q_{i,y_{i}}\|_{P_{i,y_{i}}}\right)(1-\|T\|_{\Pi}).$ (2)
If the Markov operator $Q_{i,y_{i}}$ is positive semi-definite for each $i$ and $y_{i}$, then the right-hand-side can be tightened to $1-\|T\|_{\Pi}$.
The bounds in (2) implies that if the convergence rates of all chains associated with $Q_{i,y_{i}}$ are uniformly bounded away from 1 when $i$ and $y_{i}$ vary, then the ratio of the spectral gaps of $\hat{T}$ and $T$ is in $(0,2)$. In this case, exact Gibbs and hybrid Gibbs have similar/comparable convergence rates.
We also note that (2) has the form
$\left(\inf_{i}\text{Spectral gap of chain }C_{i}\right)\times\text{Spectral gap of chain }A\leq\text{Spectral gap of chain }B.$
Interestingly, bounds of this form have appeared in various other settings that are quite distinct from the one at hand (Carlen et al., 2003; Madras and Randall, 2002; Qin, 2023). Previously, the only quantitative convergence bound for general hybrid random-scan Gibbs samplers that we know of is from Roberts and Rosenthal (1998). In the context of Example 3, Roberts and Rosenthal (1998) established a relation that bears some similarity to the first inequality in (2). However, their proof relied on the strong assumptions that the chain associated with $Q_{i,y_{i}}$ is uniformly ergodic for all $i$ and $y_{i}$, and the chain associated with $T$ is uniformly ergodic as well.
For the data augmentation sampler, we have two sets of bounds. The first is reminiscent of (2). Loosely speaking, it goes like this:
$\left(1-\sup_{y_{2}}\|Q_{2,y_{2}}\|_{P_{2,y_{2}}}\right)(1-\|S\|_{\Pi})\leq 1-\|\hat{S}\|_{\Pi}\leq\left(1+\sup_{y_{2}}\|Q_{2,y_{2}}\|_{P_{i,y_{i}}}\right)(1-\|S\|_{\Pi}).$ (3)
If the Markov operator $Q_{2,y_{2}}$ is positive semi-definite for each $y_{2}$, then the right-hand-side can be tightened to $1-\|S\|_{\Pi}$. The first inequality in (3) generalizes Theorem 6 of Jones et al. (2014). The proof of that theorem leads to a bound that is similar to the first inequality in (3), but it holds only if the chain associated with $Q_{2,y_{2}}$ is uniformly ergodic for every $y_{2}$. The second set of bounds loosely goes like this: for any positive even integer $t$,
$t(1-\|\hat{S}\|_{\Pi})\geq 1-\|\hat{S}\|_{\Pi}^{t}\geq 1-\|S\|_{\Pi}-\sup_{y_{1}}\int_{\mathsf{X}}\|Q_{2,\psi_{2}(x)}\|_{P_{2,\psi_{2}(x)}}^{t}P_{1,y_{1}}(\mathrm{d}x).$ (4)
If $Q_{2,y_{2}}$ is positive semi-definite for each $y_{2}$, then (4) hold for odd values of $t$ as well. The first inequality in (4) is a version of Bernoulli’s inequality. When $\|S\|_{\Pi}<1$, the second inequality is nontrivial when
$\lim_{t\to\infty}\sup_{y_{1}}\int_{\mathsf{X}}\|Q_{2,\psi_{2}(x)}\|_{P_{2,\psi_{2}(x)}}^{t}P_{1,y_{1}}(\mathrm{d}x)=0.$
This is a weaker condition than $\sup_{y_{2}}\|Q_{2,y_{2}}\|_{P_{2,y_{2}}}<1$, which must hold for the lower bound on $1-\|\hat{S}\|_{\Pi}$ in (3) to be non-trivial. However, in terms of bounding $1-\|\hat{S}\|_{\Pi}$ from below, (3) can be sharper than (4) if $\sup_{y_{2}}\|Q_{2,y_{2}}\|_{P_{2,y_{2}}}<1$. The lower bound on $1-\|\hat{S}\|_{\Pi}$ in (4) is a generalization of a result in Łatuszyński and Rudolf (2014), who studied a hybrid slice sampler. In Section 5.4, it is shown that this hybrid slice sampler is in fact a special case of the hybrid data augmentation sampler.
Remark 8.
We were not able to obtain an analogue of (4) for the random-scan Gibbs chain that was sufficiently simple and useful.
3 Preliminaries
This section reviews the $L^{2}$ theory for Markov chains. Let $(\mathsf{Y},\mathcal{F},\omega)$ be a probability space. The linear space $L^{2}(\omega)$ consists of functions square-integrable function $f:\mathsf{Y}\to\mathbb{R}$ such that $\int_{\mathsf{Y}}f(y)^{2}\,\omega(\mathrm{d}y)<\infty$. Two functions in $L^{2}(\omega)$ are considered identical if they are equal almost everywhere. For $f,g\in L^{2}(\omega)$ and $c\in\mathbb{R}$, $(f+g)(y)=f(y)+g(y)$, and $(cf)(y)=cf(y)$. The inner product of $f$ and $g$ is $\langle f,g\rangle_{\omega}=\int_{\mathsf{Y}}f(y)g(y)\,\omega(\mathrm{d}y)$. The norm of $f$ is $\|f\|_{\omega}=\sqrt{\langle f,f\rangle_{\omega}}$. Let $L_{0}^{2}(\omega)$ be the subspace of $L^{2}(\omega)$ consisting of functions $f$ such that $\omega f:=\int_{\mathsf{Y}}f(y)\,\omega(\mathrm{d}y)=0$. It can be shown that $(L^{2}(\omega),\langle\cdot,\cdot,\rangle_{\omega})$ and $(L_{0}^{2}(\omega),\langle\cdot,\cdot,\rangle_{\omega})$ are real Hilbert spaces. A probability measure $\mu:\mathcal{F}\to[0,1]$ is said to be in $L_{*}^{2}(\omega)$ if $\mathrm{d}\mu/\mathrm{d}\omega$ exists and is in $L^{2}(\omega)$. The $L^{2}$ distance between $\mu,\nu\in L_{*}^{2}(\omega)$ is
$\|\mu-\nu\|_{\omega}:=\sup_{f\in L_{0}^{2}(\omega)}(\mu f-\nu f)=\left\|\frac{\mathrm{d}\mu}{\mathrm{d}\omega}-\frac{\mathrm{d}\nu}{\mathrm{d}\omega}\right\|_{\omega},$
where $\mu f$ and $\nu f$ are defined analogously to $\omega f$. Let $K:\mathsf{Y}\times\mathcal{F}\to[0,1]$ be an Mtk such that
$\omega K(\cdot):=\int_{\mathsf{Y}}\omega(\mathrm{d}y)\,K(y,\cdot)=\omega(\cdot).$
For any positive integer $t$, let $K^{t}$ be the $t$-step Mtk, so that $K^{1}=K$ and $K^{t+1}(y,\cdot)=\int_{\mathsf{Y}}K(y,\mathrm{d}z)K^{t}(z,\cdot)$. Note that $K$ can be seen as a linear operator on $L_{0}^{2}(\omega)$: for $f\in L_{0}^{2}(\omega)$,
$Kf(\cdot)=\int_{\mathsf{Y}}K(\cdot,\mathrm{d}y)f(y).$
We refer to $K$ as a Markov operator. The operator corresponding to the $t$-step Mtk $K^{t}(\cdot,\cdot)$ is precisely the operator $K^{t}=\prod_{i=1}^{t}K$. The operator $K$ is self-adjoint, i.e., $\langle Kf,g\rangle_{\omega}=\langle f,Kg\rangle_{\omega}$ for $f,g\in L_{0}^{2}(\omega)$, if and only if the corresponding chain is reversible with respect to $\omega$. We say $K$ is positive semi-definite if it is self-adjoint, and $\langle f,Kf\rangle_{\omega}\geq 0$ for $f\in L_{0}^{2}(\omega)$. With a bit of work, one can show that $T$ and $S$ from Section 2 are both positive semi-definite linear operators. By Jensen’s inequality,
$\|K\|_{\omega}:=\sup_{f\in L_{0}^{2}(\omega)\setminus\{0\}}\frac{\|Kf\|_{\omega}}{\|f\|_{\omega}}\leq 1.$
Assume that $K$ is self-adjoint. Then
$\|K\|_{\omega}=\max\left\{-\inf_{f\in L_{0}^{2}(\omega)\setminus\{0\}}\frac{\langle f,Kf\rangle_{\omega}}{\|f\|_{\omega}^{2}},\sup_{f\in L_{0}^{2}(\omega)\setminus\{0\}}\frac{\langle f,Kf\rangle_{\omega}}{\|f\|_{\omega}^{2}}\right\},$ (5)
and $\|K^{t}\|=\|K\|^{t}$ for any positive integer $t$. The Dirichlet form of a function $f\in L_{0}^{2}(\omega)$ with respect to $K$ is defined to be $\|f\|_{\omega}^{2}-\langle f,Kf\rangle_{\omega}$.
Under reversibility, one can find $C:L_{*}^{2}(\omega)\to[0,\infty)$ such that
$\|\mu K^{t}-\omega\|_{\omega}\leq C(\mu)\rho^{t}$
for every $\mu\in L_{*}^{2}(\omega)$ and $t\geq 1$ if and only if $\rho\geq\|K\|_{\omega}$, where $\mu K^{t}(\cdot)=\int_{\mathsf{Y}}\mu(\mathrm{d}y)\,K^{t}(y,\cdot)$ (Roberts and Rosenthal, 1997). In this sense, $\|K\|_{\omega}$ defines the convergence rate of the corresponding reversible chain. The absolute spectral gap of $K$ is defined to be $1-\|K\|_{\omega}$.
4 Main Results
This section provides convergence rate bounds for the hybrid Gibbs samplers in the two settings introduced in Section 2. We also explain why, in the context of Example 3, the chains associated with $Q_{i,y_{i}}$ and $R_{i,y_{i}}$ have the same convergence rate.
4.1 Random-scan Gibbs chain
Recall that $T(x,A)=\sum_{i=1}^{n}p_{i}P_{i,\psi_{i}(x)}(A)$ is the Mtk of the random-scan Gibbs chain, which is reversible with respect to $\Pi$. Thus, $T$ can be regarded as a self-adjoint operator on $L_{0}^{2}(\Pi)$. In fact, by Lemma 2, it is also positive semi-definite. The convergence rate of the Gibbs chain is given by $\|T\|_{\Pi}$. For $M_{i}$-a.e. $y_{i}\in\mathsf{Y}_{i}$, the Mtk $Q_{i,y_{i}}$ is reversible with respect to $P_{i,y_{i}}$. The convergence rate of the corresponding chain is given by $\|Q_{i,y_{i}}\|_{P_{i,y_{i}}}$. The hybrid Gibbs chain of interest is associated with the Mtk $\hat{T}(x,A)=\sum_{i=1}^{n}p_{i}Q_{i,\psi_{i}(x)}(x,A)$. Recall from Lemma 4 that this chain is reversible with respect to $\Pi$, so its convergence rate is given by $\|\hat{T}\|_{\Pi}$.
Our goal is to bound $\|\hat{T}\|_{\Pi}$ in terms of $\|T\|_{\Pi}$ and an upper bound on $\|Q_{i,y_{i}}\|_{P_{i,y_{i}}}$. We first state a bound in terms of Dirichlet forms.
Theorem 9.
Suppose that there exist constants $C_{L},C_{U}\in[-1,1]$ such that
$C_{L}\leq\frac{\langle f,Q_{i,y_{i}}f\rangle_{P_{i,y_{i}}}}{\|f\|_{P_{i,y_{i}}}^{2}}\leq C_{U}$ (6)
for $i=1,\dots,n$, $M_{i}$-a.e. $y_{i}\in\mathsf{Y}_{i}$, and $f\in L_{0}^{2}(P_{i,y_{i}})\setminus\{0\}$. Then, for $f\in L_{0}^{2}(\Pi)$,
$(1-C_{L})\langle f,Tf\rangle_{\Pi}+C_{L}\|f\|_{\Pi}^{2}\leq\langle f,\hat{T}f\rangle_{\Pi}\leq(1-C_{U})\langle f,Tf\rangle_{\Pi}+C_{U}\|f\|_{\Pi}^{2}.$ (7)
In terms of Dirichlet forms, (7) is equivalent to
$(1-C_{U})(\|f\|_{\Pi}^{2}-\langle f,Tf\rangle_{\Pi})\leq\|f\|_{\Pi}^{2}-\langle f,\hat{T}f\rangle_{\Pi}\leq(1-C_{L})(\|f\|_{\Pi}^{2}-\langle f,Tf\rangle_{\Pi}).$
Proof.
Let $f\in L_{0}^{2}(\Pi)$. By linearity, (1) and the concentration of $P_{i,y_{i}}$ on $\psi^{-1}_{i}(y_{i})$,
$\langle f,\hat{T}f\rangle_{\Pi}$ $=\sum_{i=1}^{n}p_{i}\int_{\mathsf{Y}_{i}}M_{i}(\mathrm{d}y_{i})\int_{\mathsf{X}}P_{i,y_{i}}(\mathrm{d}x)f(x)\int_{\mathsf{X}}Q_{i,\psi_{i}(x)}(x,\mathrm{d}x^{\prime})f(x^{\prime})$ (8)
$=\sum_{i=1}^{n}p_{i}\int_{\mathsf{Y}_{i}}M_{i}(\mathrm{d}y_{i})\int_{\mathsf{X}}P_{i,y_{i}}(\mathrm{d}x)f(x)\int_{\mathsf{X}}Q_{i,y_{i}}(x,\mathrm{d}x^{\prime})f(x^{\prime}).$
For $i=1,\dots,n$, since, by (1),
$\|f\|_{\Pi}^{2}=\int_{\mathsf{Y}_{i}}\int_{\mathsf{X}}f(x)^{2}\,P_{i,y_{i}}(\mathrm{d}x)\,M_{i}(\mathrm{d}x)<\infty,$
it holds that $f-P_{i,y_{i}}f$ can be regarded as a function in $L_{0}^{2}(P_{i,y_{i}})$ for $M_{i}$-a.e. $y_{i}\in\mathsf{Y}_{i}$. Here, $P_{i,y_{i}}f=\int_{\mathsf{X}}P_{i,y_{i}}(\mathrm{d}x)f(x)$ can be regarded as a constant function in $L^{2}(P_{i,y_{i}})$, and the inner product between a function in $L_{0}^{2}(P_{i,y_{i}})$ and a constant function is zero. For $M_{i}$-a.e. $y_{i}\in\mathsf{Y}_{i}$, $Q_{i,y_{i}}(f-P_{i,y_{i}}f)\in L_{0}^{2}(P_{i,y_{i}})$. Combining these observations, we see
$\int_{\mathsf{X}}P_{i,y_{i}}(\mathrm{d}x)f(x)\int_{\mathsf{X}}Q_{i,y_{i}}(x,\mathrm{d}x^{\prime})f(x^{\prime})$ (9)
$=$ $\int_{\mathsf{X}}P_{i,y_{i}}(\mathrm{d}x)[f(x)-P_{i,y_{i}}f+P_{i,y_{i}}f]\int_{\mathsf{X}}Q_{i,y_{i}}(x,\mathrm{d}x^{\prime})[f(x^{\prime})-P_{i,y_{i}}f+P_{i,y_{i}}f]$
$=$ $\langle f-P_{i,y_{i}}f,Q_{i,y_{i}}(f-P_{i,y_{i}}f)\rangle_{P_{i,y_{i}}}+\langle P_{i,y_{i}}f,Q_{i,y_{i}}(f-P_{i,y_{i}}f)\rangle_{P_{i,y_{i}}}+$
$\langle f-P_{i,y_{i}}f,P_{i,y_{i}}f\rangle_{P_{i,y_{i}}}+\langle P_{i,y_{i}}f,P_{i,y_{i}}f\rangle_{P_{i,y_{i}}}$
$=$ $\langle f-P_{i,y_{i}}f,Q_{i,y_{i}}(f-P_{i,y_{i}}f)\rangle_{P_{i,y_{i}}}+\langle P_{i,y_{i}}f,P_{i,y_{i}}f\rangle_{P_{i,y_{i}}}.$
Then by (8),
$\langle f,\hat{T}f\rangle_{\Pi}=\sum_{i=1}^{n}p_{i}\int_{\mathsf{Y}_{i}}\left[\langle f-P_{i,y_{i}}f,Q_{i,y_{i}}(f-P_{i,y_{i}}f)\rangle_{P_{i,y_{i}}}+\langle P_{i,y_{i}}f,P_{i,y_{i}}f\rangle_{P_{i,y_{i}}}\right]M_{i}(\mathrm{d}y_{i}).$ (10)
For $g\in L_{0}^{2}(\Pi)$, let
$\tilde{P}_{i}g(x)=\int_{\mathsf{X}}P_{i,\psi_{i}(x)}(\mathrm{d}x^{\prime})\,g(x^{\prime}),\quad x\in\mathsf{X},$ (11)
so that $T=\sum_{i=1}^{n}p_{i}\tilde{P}_{i}$. By Lemma 2, $\tilde{P}_{i}$ as an operator on $L_{0}^{2}(\Pi)$ is idempotent and self-adjoint, i.e., an orthogonal projection, so
$\int_{\mathsf{Y}_{i}}\langle P_{i,y_{i}}f,P_{i,y_{i}}f\rangle_{P_{i,y_{i}}}M_{i}(\mathrm{d}y_{i})$ $=\int_{\mathsf{Y}_{i}}\left[\int_{\mathsf{X}}P_{i,y_{i}}(\mathrm{d}x)f(x)\right]^{2}M_{i}(\mathrm{d}y_{i})$ (12)
$=\int_{\mathsf{X}}\left[\int_{\mathsf{X}}P_{i,\psi_{i}(x)}(\mathrm{d}x)f(x)\right]^{2}\Pi(\mathrm{d}x)$
$=\langle\tilde{P}_{i}f,\tilde{P}_{i}f\rangle_{\Pi}$
$=\langle f,\tilde{P}_{i}f\rangle_{\Pi}.$
Then by (10),
$\langle f,\hat{T}f\rangle_{\Pi}=\sum_{i=1}^{n}p_{i}\int_{\mathsf{Y}_{i}}\langle f-P_{i,y_{i}}f,Q_{i,y_{i}}(f-P_{i,y_{i}}f)\rangle_{P_{i,y_{i}}}M_{i}(\mathrm{d}y_{i})+\langle f,Tf\rangle_{\Pi}.$
By (6), for each $i$ and $M_{i}$-a.e. $y_{i}\in\mathsf{Y}_{i}$
$C_{L}\|f-P_{i,y_{i}}f\|_{P_{i,y_{i}}}^{2}\leq\langle f-P_{i,y_{i}}f,Q_{i,y_{i}}(f-P_{i,y_{i}}f)\rangle_{P_{i,y_{i}}}\leq C_{U}\|f-P_{i,y_{i}}f\|_{P_{i,y_{i}}}^{2}.$
Meanwhile, by the fact that $\tilde{P}_{i}$ is an orthogonal projection,
$\int_{\mathsf{Y}_{i}}\|f-P_{i,y_{i}}f\|_{P_{i,y_{i}}}^{2}M_{i}(\mathrm{d}y_{i})$ $=\int_{\mathsf{Y}_{i}}\left[\int_{\mathsf{X}}f(x)^{2}\,P_{i,y_{i}}(\mathrm{d}x)-\left(\int_{\mathsf{X}}f(x)\,P_{i,y_{i}}(\mathrm{d}x)\right)^{2}\right]M_{i}(\mathrm{d}y_{i})$ (13)
$=\|f\|_{\Pi}^{2}-\langle\tilde{P}_{i}f,\tilde{P}_{i}f\rangle_{\Pi}$
$=\|f\|_{\Pi}^{2}-\langle f,\tilde{P}_{i}f\rangle_{\Pi}.$
Combining the three most recent displays yields (7). ∎
Based on Theorem 9, one can immediately derive the following corollary, where (14) gives the precise form of (2).
Corollary 10.
Suppose that there exists a constant $C\in[0,1]$ such that $\|Q_{i,y_{i}}\|_{P_{i,y_{i}}}\leq C$ for $i=1,\dots,n$ and $M_{i}$-a.e. $y_{i}\in\mathsf{Y}_{i}$. Then
$(1-C)(1-\|T\|_{\Pi})\leq 1-\|\hat{T}\|_{\Pi}\leq(1+C)(1-\|T\|_{\Pi}).$ (14)
If, furthermore, $Q_{i,y_{i}}$ is positive semi-definite for $i=1,\dots,n$ and $M_{i}$-a.e. $y_{i}\in\mathsf{Y}_{i}$, then
$1-\|\hat{T}\|_{\Pi}\leq 1-\|T\|_{\Pi}.$ (15)
Proof.
For $i=1,\dots,n$, define $\tilde{P}_{i}$ as in (11), so that $T=\sum_{i=1}^{n}p_{i}\tilde{P}_{i}$. Since the $\tilde{P}_{i}$’s are orthogonal projections, they are positive semi-definite. Then $T$ is positive semi-definite as well. By (5),
$0\leq\inf_{f\in L_{0}^{2}(\Pi)\setminus\{0\}}\frac{\langle f,Tf\rangle_{\Pi}}{\|f\|_{\Pi}^{2}}\leq\sup_{f\in L_{0}^{2}(\Pi)\setminus\{0\}}\frac{\langle f,Tf\rangle_{\Pi}}{\|f\|_{\Pi}^{2}}=\|T\|_{\Pi}.$
First consider the case that $Q_{i,y_{i}}$ is not necessarily always positive semi-definite. By (7) in Theorem 9, for $f\in L_{0}^{2}(\Pi)\setminus\{0\}$,
$(1+C)\frac{\langle f,Tf\rangle_{\Pi}}{\|f\|_{\Pi}^{2}}-C\leq\frac{\langle f,\hat{T}f\rangle_{\Pi}}{\|f\|_{\Pi}^{2}}\leq(1-C)\frac{\langle f,Tf\rangle_{\Pi}}{\|f\|_{\Pi}^{2}}+C.$
Then
$(1+C)\|T\|_{\Pi}-C\leq$ $\sup_{f\in L_{0}^{2}(\omega)\setminus\{0\}}\frac{\langle f,\hat{T}f\rangle_{\Pi}}{\|f\|_{\Pi}^{2}}\leq(1-C)\|T\|_{\Pi}+C,$
$-C\leq$ $\inf_{f\in L_{0}^{2}(\omega)\setminus\{0\}}\frac{\langle f,\hat{T}f\rangle_{\Pi}}{\|f\|_{\Pi}^{2}}\leq(1-C)\|T\|_{\Pi}+C.$
By (5),
$\|\hat{T}\|_{\Pi}$ $\geq\max\left\{(1+C)\|T\|_{\Pi}-C,\,-(1-C)\|T\|_{\Pi}-C\right\}=(1+C)\|T\|_{\Pi}-C,$
$\|\hat{T}\|_{\Pi}$ $\leq\max\left\{(1-C)\|T\|_{\Pi}+C,\,C\right\}=(1-C)\|T\|_{\Pi}+C.$
This gives (14). Suppose now that $Q_{i,y_{i}}$ is positive semi-definite for $i=1,\dots,n$ and $M_{i}$-a.e. $y_{i}\in\mathsf{Y}_{i}$. By (7) in Theorem 9, for $f\in L_{0}^{2}(\Pi)\setminus\{0\}$, we have
$\frac{\langle f,Tf\rangle_{\Pi}}{\|f\|_{\Pi}^{2}}\leq\frac{\langle f,\hat{T}f\rangle_{\Pi}}{\|f\|_{\Pi}^{2}}\leq(1-C)\frac{\langle f,Tf\rangle_{\Pi}}{\|f\|_{\Pi}^{2}}+C.$
Applying (5) yields $\|T\|_{\Pi}\leq\|\hat{T}\|_{\Pi}$. This gives (15). ∎
Remark 11.
The kernel $Q_{i,y_{i}}$ can be a multi-step Mtk, i.e., $Q_{i,y_{i}}(\cdot,\cdot)=\mathcal{Q}_{i,y_{i}}^{t_{i,y_{i}}}(\cdot,\cdot)$ for some positive integer $t_{i,y_{i}}$, where $\mathcal{Q}_{i,y_{i}}$ is reversible with respect to $P_{i,y_{i}}$. In this case, the bound in Corollary 10 (and in Corollaries 13 and 16 later on) holds with $\|Q_{i,y_{i}}\|_{P_{i,y_{i}}}=\|\mathcal{Q}_{i,y_{i}}^{t_{i,y_{i}}}\|_{P_{i,y_{i}}}=\|\mathcal{Q}_{i,y_{i}}\|_{P_{i,y_{i}}}^{t_{i,y_{i}}}$, where the second equality holds because the operator $\mathcal{Q}_{i,y_{i}}$ is self-adjoint.
4.2 Data augmentation chain
Recall that $S(y_{1},A)=\int_{\mathsf{X}}P_{1,y_{1}}(\mathrm{d}x)\,P_{2,\psi_{2}(x)}(\psi_{1}^{-1}(A))$ is the Mtk of a data augmentation chain. The chain is reversible with respect to $M_{1}$, and its convergence rate is determined by $\|S\|_{M_{1}}$. For $M_{2}$-a.e. $y_{2}\in\mathsf{Y}_{2}$, $Q_{2,y_{2}}$ is reversible with respect to $P_{2,y_{2}}$, and the corresponding chain’s convergence rate is $\|Q_{2,y_{2}}\|_{P_{2,y_{2}}}$. The hybrid chain of interest is associated with the Mtk $\hat{S}(y_{1},A)=\int_{\mathsf{X}}P_{1,y_{1}}(\mathrm{d}x)\,Q_{2,\psi_{2}(x)}(x,\psi_{1}^{-1}(A))$. By Lemma 7, $\hat{S}$ is reversible with respect to $M_{1}$. Our goal is to bound $\|\hat{S}\|_{M_{1}}$ in terms of $\|S\|_{M_{1}}$ and an upper bound on $\|Q_{2,y_{2}}\|_{P_{2,y_{2}}}$. We will give two sets of bounds.
4.2.1 The first bound
We first give a comparison result involving Dirichlet forms.
Theorem 12.
Suppose that there exist constants $C_{L},C_{U}\in[-1,1]$ such that
$C_{L}\leq\frac{\langle f,Q_{2,y_{2}}f\rangle_{P_{2,y_{2}}}}{\|f\|_{P_{2,y_{2}}}^{2}}\leq C_{U}$
for $M_{2}$-a.e. $y_{2}\in\mathsf{Y}_{2}$ and $f\in L_{0}^{2}(P_{2,y_{2};\psi_{1}})\setminus\{0\}$, where $L_{0}^{2}(P_{2,y_{2};\psi_{1}})$ is the set of functions $g\in L_{0}^{2}(P_{2,y_{2}})$ such that $g(x)$ depends on $x$ only through $\psi_{1}(x)$. Then, for $f\in L_{0}^{2}(M_{1})$,
$(1-C_{L})\langle f,Sf\rangle_{M_{1}}+C_{L}\|f\|_{M_{1}}^{2}\leq\langle f,\hat{S}f\rangle_{M_{1}}\leq(1-C_{U})\langle f,Sf\rangle_{M_{1}}+C_{U}\|f\|_{M_{1}}^{2},$ (16)
i.e.,
$(1-C_{U})(\|f\|_{M_{1}}^{2}-\langle f,Sf\rangle_{M_{1}})\leq\|f\|_{M_{1}}^{2}-\langle f,\hat{S}f\rangle_{M_{1}}\leq(1-C_{L})(\|f\|_{M_{1}}^{2}-\langle f,Sf\rangle_{M_{1}}).$
The proof of this theorem is similar to that of Theorem 9, and it will be relegated to Appendix B. Using Theorem 12, we can immediately establish the following corollary, which gives the precise form of (3).
Corollary 13.
Suppose that there exists a constant $C\in[0,1]$ such that $\|Q_{2,y_{2}}\|_{P_{2,y_{2}}}\leq C$ for $M_{2}$-a.e. $y_{2}\in\mathsf{Y}_{2}$. Then
$(1-C)(1-\|S\|_{M_{1}})\leq 1-\|\hat{S}\|_{M_{1}}\leq(1+C)(1-\|S\|_{M_{1}})$
If, furthermore, $Q_{2,y_{2}}$ is positive semi-definite for $M_{2}$-a.e. $y_{2}\in\mathsf{Y}_{2}$, then
$1-\|\hat{S}\|_{M_{1}}\leq 1-\|S\|_{M_{1}}.$
Proof.
Let $f\in L_{0}^{2}(M_{1})$. Then $f\circ\psi_{1}\in L_{0}^{2}(\Pi)$, where $\circ$ denotes function composition. For $i=1,2$, let $\tilde{P}_{i}:L_{0}^{2}(\Pi)\to L_{0}^{2}(\Pi)$ be defined as in (11), i.e., $\tilde{P}_{i}g(x)=\int_{\mathsf{X}}P_{i,\psi_{i}(x)}(\mathrm{d}x^{\prime})\,g(x^{\prime})$ for $x\in\mathsf{X}$. Note that $\tilde{P}_{1}(f\circ\psi_{1})=f\circ\psi_{1}$ due to the concentration of $P_{1,\psi_{1}(x)}(\cdot)$ on the set $\psi^{-1}_{1}(\psi_{1}(x))=\{x^{\prime}\in\mathsf{X}:\,\psi_{1}(x^{\prime})=\psi_{1}(x)\}$. Recall that $\tilde{P}_{1}$ and $\tilde{P}_{2}$ are orthogonal projections. Then
$\langle f,Sf\rangle_{M_{1}}=\langle f\circ\psi_{1},\tilde{P}_{1}\tilde{P}_{2}(f\circ\psi_{1})\rangle_{\Pi}=\langle\tilde{P}_{2}(f\circ\psi_{1}),\tilde{P}_{2}(f\circ\psi_{1})\rangle_{\Pi}\geq 0.$
Recall that $S$ is self-adjoint, so it is furthermore positive semi-definite. The desired result then follows from Theorem 12 through an argument that is analogous to the proof of Corollary 10. ∎
Prior to our work, the most recent result that relates $\|\hat{S}\|_{M_{1}}$ to $\|S\|_{M_{1}}$ that we know of is from Jones et al. (2014). The proof of their Theorem 6 implicitly states the following:
Proposition 14.
In the context of Example 6, suppose that there exists $C\in[0,1]$ such that, for $y_{2}\in\mathsf{Y}_{2}$, $x_{-1}\in\mathsf{Y}_{1}$, and $A\in\mathcal{F}_{1}$,
$R_{2,y_{2}}(x_{-1},A)\geq(1-C)\,\Pi_{2,y_{2}}(A).$ (17)
Then
$1-\|\hat{S}\|_{M_{1}}\geq(1-C)(1-\|S\|_{M_{1}}).$
In the context of Example 6, the inequality (17) ensures that the Markov chain associated with $Q_{2,y_{2}}$ is uniformly ergodic with convergence rate $C$ since it is equivalent to the following Doeblin’s condition:
$Q_{2,y_{2}}(x,\cdot)\geq(1-C)P_{2,y_{2}}(\cdot),\quad x\in\mathsf{X}.$
In Corollary 13, we relax this condition to $\|Q_{2,y_{2}}\|_{P_{2,y_{2}}}\leq C$, which roughly equates to the chain associated with $Q_{2,y_{2}}$ being geometrically ergodic with convergence rate $C$ (Roberts and Rosenthal, 1997; Roberts and Tweedie, 2001).
4.2.2 The second bound
The following theorem is proved in Appendix C. It takes inspiration from Łatuszyński and Rudolf (2014), who studied the convergence rates of hybrid slice samplers, which can be regarded as hybrid data augmentation samplers targeting certain uniform distributions.
Theorem 15.
Let $\gamma:\mathsf{Y}_{2}\to[0,1]$ be a measurable function such that $\|Q_{2,y_{2}}\|_{P_{2,y_{2}}}\leq\gamma(y_{2})$ for $M_{2}$-a.e. $y_{2}\in\mathsf{Y}_{2}$. Let $t$ be some positive integer, and let $\alpha_{t}\in[0,1]$ be such that
$\int_{\mathsf{X}}\gamma(\psi_{2}(x))^{t}\,P_{1,y_{1}}(\mathrm{d}x)\leq\alpha_{t}$
for $M_{1}$-a.e. $y_{1}\in\mathsf{Y}_{1}$. If either (i) $t$ is even, or (ii) $Q_{2,y_{2}}$ is positive semi-definite for $M_{2}$-a.e. $y_{2}\in\mathsf{Y}_{2}$, then, for $f\in L_{0}^{2}(M_{2})\setminus\{0\}$,
$0\leq\frac{\langle f,\hat{S}f\rangle_{M_{1}}^{t}}{\|f\|_{M_{1}}^{2t}}\leq\frac{\langle f,Sf\rangle_{M_{1}}}{\|f\|_{M_{1}}^{2}}+\alpha_{t},$ (18)
i.e.,
$1-\frac{\langle f,Sf\rangle_{M_{1}}}{\|f\|_{M_{1}}^{2}}-\alpha_{t}\leq 1-\frac{\langle f,\hat{S}f\rangle_{M_{1}}^{t}}{\|f\|_{M_{1}}^{2t}}\leq 1$
Theorem 15 gives the following bound regarding spectral gaps, which gives the precise form of (4).
Corollary 16.
For each positive integer $t$, define $\alpha_{t}$ as in Theorem 15. Then, if $t$ is even, or if $Q_{2,y_{2}}$ is positive semi-definite for $M_{2}$-a.e. $y_{2}\in\mathsf{Y}_{2}$,
$t(1-\|\hat{S}\|_{M_{1}})\geq 1-\|\hat{S}\|_{M_{1}}^{t}\geq 1-\|S\|_{M_{1}}-\alpha_{t}.$
Proof.
The first inequality is the Bernoulli inequality $t(1-x)\geq 1-x^{t}$ which holds for $x\in[-1,1]$. The second inequality follows from Theorem 15 and (5). ∎
Corollary 16 can be regarded as a direct generalization of Theorem 8 of Łatuszyński and Rudolf (2014). The latter gives a similar bound for a hybrid slice sampler. See Section 5.4 for more details.
4.3 The relationship between $Q_{i,y_{i}}$ and $R_{i,y_{i}}$
Consider Examples 1 and 3, where $P_{i,y_{i}}$ is defined in terms of $\Pi_{i,y_{i}}$, and $Q_{i,y_{i}}$ is defined in terms of $R_{i,y_{i}}$. For some researchers, it may be more natural to consider the chain associated with $R_{i,y_{i}}$ than the one associated with $Q_{i,y_{i}}$. The following proposition establishes an equivalence relationship between them in terms of $L^{2}$ norm. For $i=1,\dots,n$, and $M_{i}$-a.e. $y_{i}\in\mathsf{Y}_{i}$, $Q_{i,y_{i}}$ can be regarded as an operator on $L_{0}^{2}(P_{i,y_{i}})$, whereas $R_{i,y_{i}}$ can be seen as an operator on $L_{0}^{2}(\Pi_{i,y_{i}})$. We draw a one-to-one correspondence between the two Hilbert spaces. A linear transformation $U:L_{0}^{2}(P_{i,y_{i}})\to L_{0}^{2}(\Pi_{i,y_{i}})$ is an isometric isomorphism if $U$ is bijective and $\|Uf\|_{\Pi_{i,y_{i}}}=\|f\|_{P_{i,y_{i}}}$ for $f\in L_{0}^{2}(P_{i,y_{i}})$. The following result is proved in Appendix D.
Proposition 17.
In the context of Examples 1 and 3, for $i=1,\dots,n$ and $M_{i}$-a.e. $y_{i}\in\mathsf{Y}_{i}$, there exists an isometric isomorphism $U_{i,y_{i}}:L_{0}^{2}(P_{i,y_{i}})\to L_{0}^{2}(\Pi_{i,y_{i}})$ such that $Q_{i,y_{i}}=U_{i,y_{i}}^{-1}R_{i,y_{i}}U_{i,y_{i}}$. As a consequence, $Q_{i,y_{i}}$ is positive semi-definite if and only if $R_{i,y_{i}}$ is, and $\|Q_{i,y_{i}}\|_{P_{i,y_{i}}}=\|R_{i,y_{i}}\|_{\Pi_{i,y_{i}}}$.
5 Examples
5.1 A Metropolis-within-Gibbs sampler
Consider a special case of Examples 1 and 3. To be specific, suppose that, for $i=1,\dots,n$, $(\mathsf{X}_{i},\mathcal{B}_{i},\mu_{i})$ is $\mathbb{R}^{d_{i}}$ equipped with the Borel measure, where $d_{i}$ is a positive integer. Assume that $\Pi$ admits a density function with respect to $\prod_{i=1}^{n}\mu_{i}$ that is proportional to $x\mapsto e^{-\xi(x)}$, where $\xi:\mathsf{X}\to\mathbb{R}$ is a twice differentiable function. For $y_{i}\in\mathsf{Y}_{i}=\psi_{i}(\mathsf{X})=\mathbb{R}^{\sum_{j\neq i}d_{j}}$, define $\xi_{i,y_{i}}:\mathbb{R}^{d_{i}}\to\mathbb{R}$ as follows: $\xi_{i,y_{i}}(x_{i})=\xi(x_{1},\dots,x_{n})$ whenever $y_{i}=x_{-i}$. Then the conditional distribution $\Pi_{i,y_{i}}(\cdot)$ admits a density function with respect to $\mu_{i}$ that is proportional to $x_{i}\mapsto e^{-\xi_{i,y_{i}}(x_{i})}$. We focus on the case where $\xi_{i,y_{i}}$ is well-conditioned. To be specific, suppose that, for $i=1,\dots,n$ and $y_{i}\in\mathsf{Y}_{i}$, there exist positive numbers $r_{i,y_{i}}$ and $L_{i,y_{i}}$ such that, for $x_{i},u\in\mathbb{R}^{d_{i}}$,
$\frac{r_{i,y_{i}}}{2}\|u\|^{2}\leq\xi_{i,y_{i}}(x_{i}+u)-\xi_{i,y_{i}}(x_{i})-\langle\nabla\xi_{i,y_{i}}(x_{i}),u\rangle\leq\frac{L_{i,y_{i}}}{2}\|u\|^{2},$
[…]
$T((x_{1},\dots,x_{n}),\mathrm{d}(x_{1}^{\prime},\dots,x_{n}^{\prime}))=\sum_{i=1}^{n}p_{i}\Pi_{i,x_{-i}}(\mathrm{d}x_{i}^{\prime})\,\delta_{x_{-i}}(\mathrm{d}x_{-i}^{\prime}).$
The hybrid version has Mtk
$\hat{T}((x_{1},\dots,x_{n}),\mathrm{d}(x_{1}^{\prime},\dots,x_{n}^{\prime}))=\sum_{i=1}^{n}p_{i}R_{i,x_{-i}}(x_{i},\mathrm{d}x_{i}^{\prime})\,\delta_{x_{-i}}(\mathrm{d}x^{\prime}_{-i}).$
To be concrete, let $R_{i,y_{i}}$ be the Mtk of a random-walk Metropolis algorithm that proceeds as follows: Given the current state $x_{i}\in\mathbb{R}^{d_{i}}$, draw an element $x_{i}^{\prime}$ from the $d_{i}$-dimensional normal distribution with mean $x_{i}$ and variance $\sigma_{i,y_{i}}^{2}d_{i}^{-1}I_{d_{i}}$, where $\sigma_{i,y_{i}}>0$, and $I_{d_{i}}$ is the $d_{i}\times d_{i}$ identity matrix. The proposed state $x_{i}^{\prime}$ is then accepted or rejected using a Metropolis Hastings step (see, e.g., Andrieu et al., 2022b). The convergence properties of random-walk Metropolis algorithms have been studied extensively, especially for target distributions that are well-conditioned (Belloni and Chernozhukov, 2009; Dwivedi et al., 2019; Chen et al., 2020). In particular, there is the following result.
Proposition 18.
(Andrieu et al., 2022b, Theorem 1) For $i=1,\dots,n$ and $y_{i}\in\mathbb{R}^{d_{i}}$,
$1-\|R_{i,y_{i}}\|_{\Pi_{i,y_{i}}}\geq C_{0}\frac{r_{i,y_{i}}\sigma_{i,y_{i}}^{2}}{d_{i}}\exp\left(-2L_{i,y_{i}}\sigma_{i,y_{i}}^{2}\right),$
where $C_{0}$ is a positive universal constant.
Remark 19.
In fact, Theorem 1 of Andrieu et al. (2022b) only states that
$1-\sup_{f\in L_{0}^{2}(\Pi_{i,y_{i}})\setminus\{0\}}\frac{\langle f,R_{i,y_{i}}f\rangle_{\Pi_{i,y_{i}}}}{\|f\|_{\Pi_{i,y_{i}}}^{2}}\geq C_{0}\frac{r_{i,y_{i}}\sigma_{i,y_{i}}^{2}}{d_{i}}\exp\left(-2L_{i,y_{i}}\sigma_{i,y_{i}}^{2}\right).$
However, as shown in Lemma 3.1 of Baxendale (2005), $\langle f,R_{i,y_{i}}f\rangle_{\Pi_{i,y_{i}}}\geq 0$ for $f\in L_{0}^{2}(\Pi_{i,y_{i}})$. Thus, by (5),
$1-\|R_{i,y_{i}}\|_{\Pi_{i,y_{i}}}=1-\sup_{f\in L_{0}^{2}(\Pi_{i,y_{i}})\setminus\{0\}}\frac{\langle f,R_{i,y_{i}}f\rangle_{\Pi_{i,y_{i}}}}{\|f\|_{\Pi_{i,y_{i}}}^{2}}.$
Applying Corollary 10 along with Proposition 17 gives the following:
Proposition 20.
$1-\|\hat{T}\|_{\Pi}\geq C_{0}\min_{i}\inf_{y_{i}\in\mathsf{Y}_{i}}\left[\frac{r_{i,y_{i}}\sigma_{i,y_{i}}^{2}}{d_{i}}\exp\left(-2L_{i,y_{i}}\sigma_{i,y_{i}}^{2}\right)\right](1-\|T\|_{\Pi}),$
where $C_{0}$ is a universal constant. In particular, if $\sigma_{i,y_{i}}^{2}=1/(2L_{i,y_{i}})$ for each value of $(i,y_{i})$, then
$1-\|\hat{T}\|_{\Pi}\geq\frac{C_{0}}{2e}\min_{i}\inf_{y_{i}\in\mathsf{Y}_{i}}\left(\frac{1}{\kappa_{i,y_{i}}d_{i}}\right)(1-\|T\|_{\Pi}),$
where $\kappa_{i,y_{i}}=L_{i,y_{i}}/r_{i,y_{i}}$ is called the condition number of $\Pi_{i,y_{i}}$.
If there exist $\kappa<\infty$ and $d<\infty$ such that $L_{i,y_{i}}/r_{i,y_{i}}\leq\kappa$ and $d_{i}\leq d$ for each value of $(i,y_{i})$, then $(1-\|T\|_{\Pi})/(1-\|\hat{T}\|_{\Pi})=O(\kappa d)$ as $\kappa,d\to\infty$. Recall our discussion in Section 1: $(1-\|T\|_{\Pi})/(1-\|\hat{T}\|_{\Pi})$ is roughly the ratio of the mixing times of the hybrid and exact samplers. If the average per-iteration computational cost of the exact Gibbs sampler is much larger than $\kappa d$ times that of the hybrid sampler, then heuristically one may say that the hybrid sampler is more efficient than the exact one.
5.2 A proximal sampler
Consider a special case of Examples 5 and 6. To be specific, suppose that $(\mathsf{Y}_{1},\mathcal{F}_{1},\mu_{1})$ is the $d$-dimensional Euclidean space endowed with the Borel measure, where $d$ is a positive integer. Suppose that $M_{1}$, the distribution of interest, admits a density function with respect to $\mu_{1}$ that is proportional to $y_{1}\mapsto e^{-\xi(y_{1})}$, where $\xi:\mathsf{Y}_{1}\to\mathbb{R}$ is twice differentiable. Assume that there exist constants $r\in\mathbb{R}$ and $L\in[0,\infty)$ such that, for $y_{1},u\in\mathbb{R}^{d}$,
$\frac{r}{2}\|u\|^{2}\leq\xi(y_{1}+u)-\xi(y_{1})-\langle\nabla\xi(y_{1}),u\rangle\leq\frac{L}{2}\|u\|^{2}.$
Now, let $(\mathsf{Y}_{2},\mathcal{F}_{2},\mu_{2})=(\mathsf{Y}_{1},\mathcal{F}_{1},\mu_{1})$. For $y_{1}\in\mathbb{R}^{d}$, let $\Pi_{1,y_{1}}$ be a Gaussian distribution whose density is
$y_{2}\mapsto(2\pi\eta)^{-d/2}\exp\left(-\frac{\|y_{2}-y_{1}\|^{2}}{2\eta}\right),$
where $\eta$ is a positive constant. Then the joint distribution $\Pi$ has a density proportional to
$(y_{2},y_{1})\mapsto\exp\left[-\xi(y_{1})-\frac{\|y_{2}-y_{1}\|^{2}}{2\eta}\right]$
with respect to $\mu_{2}\times\mu_{1}$. The measure $M_{2}$ has a density that is proportional to the above formula with $y_{1}$ integrated out. For $M_{2}$-a.e. $y_{2}\in\mathbb{R}^{d}$, $\Pi_{2,y_{2}}$ has density function
$y_{1}\mapsto\left\{\int_{\mathbb{R}^{d}}\exp\left[-\xi(y_{1}^{\prime})-\frac{\|y_{2}-y_{1}^{\prime}\|^{2}}{2\eta}\right]\,\mathrm{d}y_{1}^{\prime}\right\}^{-1}\exp\left[-\xi(y_{1})-\frac{\|y_{2}-y_{1}\|^{2}}{2\eta}\right].$
In this case, the data augmentation chain associated with the Mtk
$S(y_{1},A)=\int_{\mathbb{R}^{d}}\Pi_{1,y_{1}}(\mathrm{d}y_{2})\,\Pi_{2,y_{2}}(A)$
is called a proximal sampler (Lee et al., 2021a). It can be particularly useful when $M_{1}$ is ill-conditioned, although later on in our analysis of its hybrid version, we focus on the case where $M_{1}$ is well-conditioned for tractability. This sampler was studied by Lee et al. (2021b) and Chen et al. (2022). In particular, using a result in Lee et al. (2021b), one can establish the following result under strong convexity.
Proposition 21.
Suppose that $r>0$. Then $\|S\|_{M_{1}}\leq 1/(1+r\eta)$.
Proof.
In the proof of Lemma 2 of Lee et al. (2021b), it was shown that, for $y_{1},y_{1}^{\prime}\in\mathbb{R}^{d}$, the 2-Wasserstein distance between $S(y_{1},\cdot)$ and $S(y_{1}^{\prime},\cdot)$ is upper bounded by $\|y_{1}-y_{1}^{\prime}\|/(1+\eta r)$. Then the 1-Wasserstein distance between the two is also subject to this upper bound. Moreover,
$\frac{1}{2}\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{2d}}\|y_{1}^{\prime}-y_{1}^{\prime\prime}\|^{2}\,S(y_{1},\mathrm{d}y_{1}^{\prime})\,S(y_{1},\mathrm{d}y_{1}^{\prime\prime})\,M_{1}(\mathrm{d}y_{1})\leq 2\int_{\mathbb{R}^{d}}\|y_{1}\|^{2}\,M_{1}(\mathrm{d}y_{1})<\infty.$
The desired result then follows from Lemma 30 of Ollivier (2009). ∎
In practice, the conditional distribution $\Pi_{2,y_{2}}(\cdot)$ may be difficult to sample from exactly. Let us now consider the hybrid version of the data augmentation chain whose Mtk has the form
$\hat{S}(y_{1},A)=\int_{\mathbb{R}^{d}}\Pi_{1,y_{1}}(\mathrm{d}y_{2})\,R_{2,y_{2}}(y_{1},A)=(2\pi\eta)^{-d/2}\int_{\mathbb{R}^{d}}\exp\left(-\frac{\|y_{2}-y_{1}\|^{2}}{2\eta}\right)R_{2,y_{2}}(y_{1},A).$
The approximating kernel $R_{2,y_{2}}$ may correspond to a number of algorithms. For illustration, we assume that $R_{2,y_{2}}$ is the Mtk of a random-walk Metropolis algorithm analogous to the one described in Section 5.1. In particular, the new state is proposed via a $d$-dimensional normal distribution whose mean is the current state, and whose variance is $\sigma^{2}d^{-1}I_{d}$ for some constant $\sigma>0$. The following lemma follows directly from Proposition 18.
Lemma 22.
Suppose that $r+1/\eta>0$. Then, for $y_{2}\in\mathbb{R}^{d}$,
$1-\|R_{2,y_{2}}\|_{\Pi_{2,y_{2}}}\geq C_{0}\frac{(r\eta+1)\sigma^{2}}{d\eta}\exp\left[-2\left(L+\frac{1}{\eta}\right)\sigma^{2}\right],$
where $C_{0}$ is a positive universal constant.
Corollaries 13 and 16, along with Proposition 17, provide bounds on $1-\|\hat{S}\|_{M_{1}}$ in terms of $1-\|S\|_{M_{1}}$ and $\|Q_{2,y_{2}}\|_{P_{2,y_{2}}}=\|R_{2,y_{2}}\|_{\Pi_{2,y_{2}}}$. In particular, combining the bound in Corollary 13 with Lemma 22 yields the following.
Proposition 23.
Suppose that $r+1/\eta>0$. Then
$1-\|\hat{S}\|_{M_{1}}\geq C_{0}\frac{(r\eta+1)\sigma^{2}}{d\eta}\exp\left[-2\left(L+\frac{1}{\eta}\right)\sigma^{2}\right](1-\|S\|_{M_{1}}),$
where $C_{0}$ is a positive universal constant.
Combining Proposition 21 with Proposition 23, we immediately have:
Proposition 24.
If $r>0$, then
$1-\|\hat{S}\|_{M_{1}}\geq C_{0}\frac{r\sigma^{2}}{d}\exp\left[-2\left(L+\frac{1}{\eta}\right)\sigma^{2}\right].$
In particular, one can choose $\eta=1/L$ and $\sigma^{2}=1/[2(L+1/\eta)]$ so that
$1-\|\hat{S}\|_{M_{1}}\geq\frac{C_{0}}{4e\kappa d},$
where $\kappa=L/r$ is the condition number of $M_{1}$.
Proposition 24 demonstrates that the hybrid data augmentation sampler achieves a spectral gap that is no smaller than a constant multiple of $(\kappa d)^{-1}$. We can readily extend these bounds to analyze the mixing time of the underlying Markov chain using conventional techniques (see, e.g., Roberts and Rosenthal, 1997, Theorem 2.1). Recall that an initial distribution $\pi_{0}$ is said to be $\beta$-warm with respect to the stationary distribution $M_{1}$ if $\pi_{0}(A)/M_{1}(A)\leq\beta$ for every measurable $A$ (Dwivedi et al., 2019). Assuming the hybrid sampler begins with a $\beta$-warm initialization, we can deduce from Theorem 2.1 of Roberts and Rosenthal (1997) that its $\epsilon$-mixing time, measured in both the total variation and $L^{2}$ distance, is on the order of $O(\kappa d\log\left(\beta/\epsilon\right))$ iterations. This exhibits the same order of magnitude as the mixing time bound for the Metropolis adjusted Langevin Algorithm (MALA) in Dwivedi et al. (2019).
A frequently adopted initialization choice is $\pi_{0}=N_{d}(y^{\star},L^{-1}I_{d})$, where $y^{\star}$ is the unique maximizer of $y_{1}\mapsto e^{-\xi(y_{1})}$ (often obtainable through optimization methods with minimal computational overhead). This particular $\pi_{0}$ corresponds to $\beta=O(\kappa^{d/2})$, resulting in a computational cost of $O\left(\kappa d(d\log(\kappa)+\log(\epsilon^{-1}))\right)$.
The above cost is of the same order as a method proposed by Chen et al. (2022), which implements the exact proximal sampler with $\eta\approx 1/(Ld)$ through a meticulous rejection sampling technique to draw from $\Pi_{2,y_{2}}(\cdot)$. (See page 7 of that work.) As noted in Chen et al. (2022), this cost is in some sense state-of-the-art. It is worth mentioning that the hybrid sampler herein may be easier to implement, as the method in Chen et al. (2022) requires accurately solving a $d$-dimensional convex optimization problem in every iteration.
5.3 Random-scan Gibbs samplers with blocked updates
For $i=1,\dots,k$ where $k$ is a positive integer, let $(\mathsf{X}_{i},\mathcal{B}_{i})$ be a Polish measurable space. Set $(\mathsf{X},\mathcal{B})=\prod_{i=1}^{k}(\mathsf{X}_{i},\mathcal{B}_{i})$, and let $\Pi$ be a probability distribution on $(\mathsf{X},\mathcal{B})$. Let $X=(X_{1},\dots,X_{k})\sim\Pi$.
For a positive integer $\ell\in[k-1]:=\{1,\dots,k-1\}$, a uniform random-scan Gibbs sampler with block size $\ell$ targeting $\Pi$ proceeds as follows: Given the current state $x$, one uniformly and randomly selects $\ell$ distinct integers from $[k]$, and call the collection of these integers $\Lambda$. The next state is then drawn from the conditional distribution of $X$ given $\psi_{\Lambda}(X)=\psi_{\Lambda}(x)$, where, for $x^{\prime}=(x^{\prime}_{1},\dots,x^{\prime}_{k})\in\mathsf{X}$, $\psi_{\Lambda}(x^{\prime})=(x^{\prime}_{i})_{i\not\in\Lambda}$, e.g., when $k=4$,
$\psi_{\{1,3\}}(x^{\prime})=\psi_{\{3,1\}}(x^{\prime})=(x^{\prime}_{2},x^{\prime}_{4}).$
In other words, this sampler randomly chooses $\ell$ coordinates to update using its conditional distribution given all other coordinates. Formally, for $\Lambda\subset[k]$, let $\mathsf{Y}_{\Lambda}=\psi_{\Lambda}(\mathsf{X})=\prod_{i\not\in\Lambda}\mathsf{X}_{i}$, $M_{\Lambda}=\Pi\circ\psi_{\Lambda}^{-1}$, and let $(P_{\Lambda,y_{\Lambda}})_{y_{\Lambda}\in\mathsf{Y}_{\Lambda}}$ be a disintegration of $\Pi$ with respect to $\psi_{\Lambda}$. That is, $M_{\Lambda}$ is the distribution of $\psi_{\Lambda}(X)=(X_{i})_{i\not\in\Lambda}$, and $P_{\Lambda,y_{\Lambda}}$ is the conditional distribution of $X$ given $\psi_{\Lambda}(X)=y_{\Lambda}$. The Mtk of the Gibbs chain is then
$T_{\ell}(x,A)=\frac{1}{{k\choose\ell}}\sum_{\Lambda\subset[k],\,|\Lambda|=\ell}P_{\Lambda,\psi_{\Lambda}(x)}(A),\quad x\in\mathsf{X},\;A\in\mathcal{B}.$
This is a special case of the random-scan Gibbs chain defined in Section 2. If we label the subsets of $[k]$ that are of cardinality $\ell$ using positive integers, then we recover the notations in Section 2. Let $\ell$ and $\ell^{\prime}$ be two block sizes such that $\ell>\ell^{\prime}$. One may then ask: is there a quantitative relationship between $\|T_{\ell}\|_{\Pi}$ and $\|T_{\ell^{\prime}}\|_{\Pi}$? This problem was considered by Qin and Wang (2023) for the case $\ell^{\prime}=\ell-1$ under the assumption that $\Pi$ admits a density function with respect to some product measure on $(\mathsf{X},\mathcal{B})$. What follows will be a direct generalization of one of their main results. We will show that the chain associated with $T_{\ell^{\prime}}$ can be viewed as a hybrid version of the Gibbs chain associated with $T_{\ell}$. Thus, one can connect $\|T_{\ell}\|_{\Pi}$ and $\|T_{\ell^{\prime}}\|_{\Pi}$ using Corollary 10 by taking $T=T_{\ell}$ and $\hat{T}=T_{\ell^{\prime}}$. Our goal is now to show that, when $\ell>\ell^{\prime}$, we can write
$T_{\ell^{\prime}}(x,\cdot)$ $:=\frac{1}{{k\choose\ell^{\prime}}}\sum_{\Lambda^{\prime}\subset[k],\,|\Lambda^{\prime}|=\ell^{\prime}}P_{\Lambda^{\prime},\psi_{\Lambda^{\prime}}(x)}(\cdot)$ (19)
$=\frac{1}{{k\choose\ell}}\sum_{\Lambda\subset[k],\,|\Lambda|=\ell}Q_{\Lambda,\psi_{\Lambda}(x)}(x,\cdot)$
for $\Pi$-a.e. $x\in\mathsf{X}$, where $Q_{\Lambda,y_{\Lambda}}$ is an Mtk that is reversible with respect to $P_{\Lambda,y_{\Lambda}}$ for $y_{\Lambda}\in\mathsf{Y}_{\Lambda}$. Interestingly, this holds when $Q_{\Lambda,y_{\Lambda}}$ also corresponds to a random-scan Gibbs sampler. Loosely speaking, we can design the following Gibbs sampler targeting the conditional distribution of $X$ given $\psi_{\Lambda}(X)=y_{\Lambda}$ that updates $\ell^{\prime}$ coordinates at a time: Given the current state $x\in\mathsf{X}$ where $\psi_{\Lambda}(x)=y_{\Lambda}$, uniformly randomly select a subset $\Lambda^{\prime}$ from the subsets of $\Lambda$ that are of cardinality $\ell^{\prime}$. The next state is then drawn from the conditional distribution of $X^{\prime}$ given $\psi_{\Lambda^{\prime}}(X^{\prime})=\psi_{\Lambda^{\prime}}(x)$, where $X^{\prime}$ is distributed as $P_{\Lambda,y_{\Lambda}}$. Formally, let $(P_{\Lambda,y_{\Lambda};\Lambda^{\prime},z_{\Lambda^{\prime}}})_{z_{\Lambda^{\prime}}\in\mathsf{Y}_{\Lambda^{\prime}}}$ be a disintegration of $P_{\Lambda,y_{\Lambda}}$ with respect to $\psi_{\Lambda^{\prime}}$. Define a Gibbs sampler targeting $P_{\Lambda,y_{\Lambda}}$ with the Mtk
$Q_{\Lambda,y_{\Lambda}}(x,A)=\frac{1}{{\ell\choose\ell^{\prime}}}\sum_{\Lambda^{\prime}\subset\Lambda,\,|\Lambda^{\prime}|=\ell^{\prime}}P_{\Lambda,y_{\Lambda};\,\Lambda^{\prime},\psi_{\Lambda^{\prime}}(x)}(A),\quad x\in\mathsf{X},\;A\in\mathcal{B}.$
By Lemma 2, this Mtk is indeed reversible with respect to $P_{\Lambda,y_{\Lambda}}$, and the corresponding operator is positive semi-definite. The following regularity condition shall be assumed regarding $(P_{\Lambda,y_{\Lambda};\Lambda^{\prime},z_{\Lambda^{\prime}}})_{z_{\Lambda^{\prime}}\in\mathsf{Y}_{\Lambda^{\prime}}}$:
  1. (A)
    Given any choice of $\Lambda$ and $\Lambda^{\prime}$, for $A\in\mathcal{B}$, the function $(y_{\Lambda},z_{\Lambda^{\prime}})\mapsto P_{\Lambda,y_{\Lambda};\,\Lambda^{\prime},z_{\Lambda^{\prime}}}(A)$ is measurable.
The following result is proved in Appendix E.
Lemma 25.
Suppose that (A) holds. Let $\Lambda\subset[k]$ and $\Lambda^{\prime}\subset\Lambda$ be such that |Λ|=normal-$|\Lambda|=\ell$ and |Λ|=superscriptnormal-$|\Lambda^{\prime}|=\ell^{\prime}$. Then $P_{\Lambda,\psi_{\Lambda}(x);\,\Lambda^{\prime},\psi_{\Lambda^{\prime}}(x)}(\cdot)=P_{\Lambda^{\prime},\psi_{\Lambda^{\prime}}(x)}(\cdot)$ for $\Pi$-a.e. $x\in\mathsf{X}$.
It follows from Lemma 25 that (19) holds for $\Pi$-a.e. $x\in\mathsf{X}$. Applying Corollary 10 shows the following:
Proposition 26.
If there exists a constant $C\in[0,1]$ such that $\|Q_{\Lambda,y_{\Lambda}}\|_{P_{\Lambda,y_{\Lambda}}}\leq C$, then
$(1-C)(1-\|T_{\ell}\|_{\Pi})\leq 1-\|T_{\ell^{\prime}}\|_{\Pi}\leq 1-\|T_{\ell}\|_{\Pi}.$
5.4 A hybrid slice sampler
Consider a special case of Examples 5 and 6. To be specific, suppose that $M_{1}$, the distribution of interest, admits a positive measurable density function with respect to some underlying $\sigma$-finite measure $\mu_{1}$ on a Polish measurable space $(\mathsf{Y}_{1},\mathcal{F}_{1})$. Assume that this density function is proportional to $\tilde{m}_{1}:\mathsf{Y}_{1}\to(0,\infty)$. Suppose further that $(\mathsf{Y}_{2},\mathcal{F}_{2},\mu_{2})$ is $[0,\infty)$ endowed with the Borel measure, and that, for $y_{1}\in\mathsf{Y}_{1}$, $\Pi_{1,y_{1}}$ is the uniform distribution on $[0,\tilde{m}_{1}(y_{1})]$. Then $\Pi$ is the uniform distribution with respect to $\mu_{2}\times\mu_{1}$ on the set
$G=\{(y_{2},y_{1})\in\mathsf{X}:\,y_{2}\in[0,\tilde{m}_{1}(y_{1})]\}.$
(It can be checked that $G$ is measurable.) The measure $M_{2}$ has a density with respect to $\mu_{2}$ given by
$m_{2}(y_{2})=\frac{\int_{\mathsf{Y}_{1}}1_{G}(y_{2},y_{1})\,\mu_{1}(\mathrm{d}y_{1})}{\int_{\mathsf{Y}_{2}}\int_{\mathsf{Y}_{1}}1_{G}(y_{2}^{\prime},y_{1})\,\mu_{1}(\mathrm{d}y_{1})\,\mu_{2}(\mathrm{d}y_{2}^{\prime})}=\frac{\int_{\mathsf{Y}_{1}}1_{G}(y_{2},y_{1})\,\mu_{1}(\mathrm{d}y_{1})}{\int_{\mathsf{Y}_{1}}\tilde{m}_{1}(y_{1})\,\mu_{1}(\mathrm{d}y_{1})}.$
For $y_{2}\in[0,\infty)$ outside an $M_{2}$ measure zero set, $\Pi_{2,y_{2}}$ is the uniform distribution on
$G_{y_{2}}=\{y_{1}\in\mathsf{Y}_{1}:\,(y_{2},y_{1})\in G\}=\{y_{1}\in\mathsf{Y}_{1}:\,y_{2}\leq\tilde{m}_{1}(y_{1})\}.$
In this scenario, the data augmentation algorithm associated with the Mtk $S(\cdot,\cdot)$ is called a slice sampler. To simulate the underlying chain, given the current state $y_{1}\in\mathsf{Y}_{1}$, one draws $y_{2}$ from the uniform distribution on $[0,\tilde{m}_{1}(y_{1})]$, then draws the next state from the uniform distribution on $G_{y_{2}}$. Regarding the spectral gap of slice samplers, the following result was recently obtained by Natarovskii et al. (2021). To state the result, for $y_{2}\in[0,\infty)$, let $v(y_{2})=\int_{\mathsf{Y}_{1}}1_{G}(y_{2},y_{1})\,\mu_{1}(\mathrm{d}y_{1})$.
Proposition 27.
(Natarovskii et al., 2021, Theorem 3.10, Proposition 3.13) Suppose that $v$ is strictly decreasing and continuously differentiable on the interval $S_{v}=(0,\sup\{y_{2}\in(0,\infty):\,v(y_{2})>0\})$. Suppose further that, for some positive integer $k$, the function $y_{2}\mapsto y_{2}v^{\prime}(y_{2})/v(y_{2})^{1-1/k}$ is decreasing on $S_{v}$. Then
$1-\|S\|_{M_{1}}\geq\frac{1}{k+1}.$
Sampling from the uniform distribution on $G_{y_{2}}$ may be difficult in practice. Instead, one can simulate one iteration of chain associated with an Mtk $R_{2,y_{2}}$, where $R_{2,y_{2}}$ is reversible with respect to $\Pi_{2,y_{2}}$. To ensure (H), we will assume that, for $A\in\mathcal{F}_{1}$, the function $(y_{2},y_{1})\mapsto R_{2,y_{2}}(y_{1},A)$ is measurable. Then we may define a hybrid chain with Mtk
$\hat{S}(y_{1},A)=\int_{\mathsf{Y}_{2}}\Pi_{1,y_{1}}(\mathrm{d}y_{2})\,R_{2,y_{2}}(y_{1},A)=\frac{1}{\tilde{m}_{1}(y_{1})}\int_{0}^{\tilde{m}(y_{1})}R_{2,y_{2}}(y_{1},A)\,\mu_{2}(\mathrm{d}y_{2}).$
Corollaries 13 and 16, along with Proposition 17, provide bounds on $1-\|\hat{S}\|_{M_{1}}$ in terms of $1-\|S\|_{M_{1}}$ and $\|Q_{2,y_{2}}\|_{P_{2,y_{2}}}=\|R_{2,y_{2}}\|_{\Pi_{2,y_{2}}}$. We shall use the upper bound from Corollary 13 and the lower bound from Corollary 16.
Proposition 28.
Let $t$ be a positive integer. If $t$ is even, or $R_{2,y_{2}}$ is positive semi-definite for $M_{2}$-a.e. $y_{2}\in\mathsf{Y}_{2}$, then
$\frac{1-\|S\|_{M_{1}}-\alpha_{t}}{t}\leq 1-\|\hat{S}\|_{M_{1}}\leq 1-\|S\|_{M_{1}},$ (20)
where $\alpha_{t}\in[0,1]$ satisfies
$\int_{0}^{\tilde{m}_{1}(y_{1})}\frac{\gamma(y_{2})^{t}}{\tilde{m}_{1}(y_{1})}\,\mu_{2}(\mathrm{d}y_{2})\leq\alpha_{t}\;\;\text{for }M_{1}\text{-a.e. }y_{1},\quad\|R_{2,y_{2}}\|_{\Pi_{2,y_{2}}}\leq\gamma(y_{2})\;\;\text{for }M_{2}\text{-a.e. }y_{2}.$
The above display recovers the main technical result of Łatuszyński and Rudolf (2014). When $R_{2,y_{2}}$ is almost surely positive semi-definite, Theorem 8 of Łatuszyński and Rudolf (2014) states the following: For each positive integer $t$,
$\frac{1-\|S\|_{M_{1}}-\beta_{t}}{t}\leq 1-\|\hat{S}\|_{M_{1}}\leq 1-\|S\|_{M_{1}},$ (21)
where $\beta_{t}\in[0,1]$ satisfies
$\left[\int_{0}^{\tilde{m}_{1}(y_{1})}\frac{\gamma(y_{2})^{2t}}{\tilde{m}_{1}(y_{1})}\,\mu_{2}(\mathrm{d}y_{2})\right]^{1/2}\leq\beta_{t}\;\;\text{for }M_{1}\text{-a.e. }y_{1},\quad\|R_{2,y_{2}}\|_{\Pi_{2,y_{2}}}\leq\gamma(y_{2})\;\;\text{for }M_{2}\text{-a.e. }y_{2}.$
By Jensen’s inequality, $\alpha_{t}\leq\beta_{t}$, so the lower bound in (20) is slightly sharper than that in (21). The reader can refer to Łatuszyński and Rudolf (2014) for some choices of practically useful $R_{2,y_{2}}$ with expressions for $\beta_{t}$.
6 Discussion
To obtain a non-trivial lower bound on $1-\|\hat{T}\|_{\Pi}$ using Corollary 10, it is imperative that $\|Q_{i,y_{i}}\|_{P_{i,y_{i}}}\leq C$ for all values of $i$ and $y_{i}$, where $C$ is some constant that is strictly less than one. Relaxing this constraint emerges as a clear avenue for prospective research. The comparative outcomes in Corollaries 10, 13, and 16 are useful only if the exact Gibbs sampler in question admits a non-vanishing spectral gap. However, even in the absence of a positive spectral gap, the comparison results based on Dirichlet forms, as expounded in Theorems 9, 12, and 15, may lead to valuable insights into the convergence behavior of the hybrid samplers (Andrieu et al., 2022a). Additionally, exploring the comparison in terms of the conductance (Lawler and Sokal, 1988), the $s$-conductance (Lovász and Simonovits, 1993), or the approximate spectral gap (Atchadé, 2021) could yield further interesting results. Acknowledgment: The first author was partially supported by the National Science Foundation through grant DMS-2112887. The third author was partially supported by the National Science Foundation through grant DMS-2210849 and an Adobe Data Science Research Award. The authors gratefully acknowledge the 2023 IRSA Conference at the University of Minnesota for catalyzing the collaborative efforts among them. Appendix
Appendix A Some Technical Lemmas
Lemma 29.
Fix $i\in\{1,\dots,n\}$. For $\Pi$-a.e. $x\in\mathsf{X}$, the probability measure $Q_{i,\psi_{i}(x)}(x,\cdot)$ is concentrated on $\psi^{-1}_{i}(\psi_{i}(x))$.
Proof.
For $M_{i}$-a.e. $y_{i}\in\mathsf{Y}_{i}$, the measure $P_{i,y_{i}}(\cdot)$ is concentrated on $\psi^{-1}_{i}(y_{i})$ and $P_{i,y_{i}}(\cdot)$ is a stationary distribution of $Q_{i,y_{i}}(\cdot,\cdot)$. Then, for such values of $y_{i}$,
$\int_{\mathsf{X}}P_{i,y_{i}}(\mathrm{d}x)\,Q_{i,y_{i}}(x,\mathsf{X}\setminus\psi^{-1}_{i}(y_{i}))=P_{i,y_{i}}(\mathsf{X}\setminus\psi^{-1}_{i}(y_{i}))=0.$
By this, (1), and the concentration of $P_{i,y_{i}}$ on $\psi^{-1}_{i}(y_{i})$,
$\int_{\mathsf{X}}\Pi(\mathrm{d}x)\,Q_{i,\psi_{i}(x)}(x,\mathsf{X}\setminus\psi^{-1}_{i}(\psi_{i}(x)))$
$=$ $\int_{\mathsf{Y}_{i}}M_{i}(\mathrm{d}y_{i})\int_{\mathsf{X}}P_{i,y_{i}}(\mathrm{d}x)\,Q_{i,\psi_{i}(x)}(x,\mathsf{X}\setminus\psi^{-1}_{i}(\psi_{i}(x)))$
$=$ $\int_{\mathsf{Y}_{i}}M_{i}(\mathrm{d}y_{i})\int_{\mathsf{X}}P_{i,y_{i}}(\mathrm{d}x)\,Q_{i,y_{i}}(x,\mathsf{X}\setminus\psi^{-1}_{i}(y_{i}))$
$=$ $0$
Note that the above integrals are well-defined because the function
[…]
where $A_{i}=\{(y_{i},x^{\prime})\in\mathsf{Y}_{i}\times\mathsf{X}:\,y_{i}\neq\psi_{i}(x^{\prime})\}$, is measurable by (H). (See, e.g., Chapter 4, Theorem 20 of Pollard, 2002) It follows that for $\Pi$-a.e. $x\in\mathsf{X}$, $Q_{i,\psi_{i}(x)}(x,\mathsf{X}\setminus\psi^{-1}_{i}(\psi_{i}(x)))=0$. ∎
Lemma 30.
[…]
$0\leq\left(\frac{\langle f,Kf\rangle}{\|f\|^{2}}\right)^{t}\leq\frac{\langle f,K^{t}f\rangle}{\|f\|^{2}}.$
If $K$ is positive semi-definite, then the inequality also for odd positive integers.
Proof.
Fix $f\in H\setminus\{0\}$. Denote by $E_{K}(\cdot)$ the spectral measure of $K$ (see, e.g., Arveson et al., 2002, Section 2.7). Note that $\langle f,E_{K}(\cdot)f\rangle$ is a measure on $\mathbb{R}$, and $\langle f,E_{K}(\mathbb{R})f\rangle=\|f\|^{2}$. Thus,
$\nu_{f}(\cdot)=\frac{\langle f,E_{K}(\cdot)f\rangle}{\|f\|^{2}}$
is a probability measure on $\mathbb{R}$. Assume that $t$ is even or $K$ is positive semi-definite. Then the function $\lambda\mapsto\lambda^{t}$ is convex on the support of $\nu_{f}$. By the spectral decomposition of $K$ and Jensen’s inequality,
$\left(\frac{\langle f,Kf\rangle}{\|f\|^{2}}\right)^{t}$ $=\left(\frac{\int_{\mathbb{R}}\lambda\langle f,E_{K}(\mathrm{d}\lambda)f\rangle}{\|f\|^{2}}\right)^{t}$
$=\left(\int_{\mathbb{R}}\lambda\nu_{f}(\mathrm{d}\lambda)\right)^{t}$
$\leq\int_{\mathbb{R}}\lambda^{t}\nu_{f}(\mathrm{d}\lambda)$
$=\frac{\langle f,K^{t}f\rangle}{\|f\|^{2}}.$
Appendix B Proof of Theorem 12
Recall the statement of Theorem 12: Theorem 12.Suppose that there exist constants $C_{L},C_{U}\in[-1,1]$ such that
$C_{L}\leq\frac{\langle f,Q_{2,y_{2}}f\rangle_{P_{2,y_{2}}}}{\|f\|_{P_{2,y_{2}}}^{2}}\leq C_{U}$
for $M_{2}$-a.e. $y_{2}\in\mathsf{Y}_{2}$ and $f\in L_{0}^{2}(P_{2,y_{2};\psi_{1}})\setminus\{0\}$, where $L_{0}^{2}(P_{2,y_{2};\psi_{1}})$ is the set of functions $g\in L_{0}^{2}(P_{2,y_{2}})$ such that $g(x)$ depends on $x$ only through $\psi_{1}(x)$. Then, for $f\in L_{0}^{2}(M_{1})$,
$(1-C_{L})\langle f,Sf\rangle_{M_{1}}+C_{L}\|f\|_{M_{1}}^{2}\leq\langle f,\hat{S}f\rangle_{M_{1}}\leq(1-C_{U})\langle f,Sf\rangle_{M_{1}}+C_{U}\|f\|_{M_{1}}^{2},$ (16)
i.e.,
$(1-C_{U})(\|f\|_{M_{1}}^{2}-\langle f,Sf\rangle_{M_{1}})\leq\|f\|_{M_{1}}^{2}-\langle f,\hat{S}f\rangle_{M_{1}}\leq(1-C_{L})(\|f\|_{M_{1}}^{2}-\langle f,Sf\rangle_{M_{1}}).$
Proof.
Let $f\in L_{0}^{2}(M_{1})$. Then $f\circ\psi_{1}\in L_{0}^{2}(\Pi)$, and $\|f\|_{M_{1}}=\|f\circ\psi_{1}\|_{\Pi}$. By (1), Lemma 2, and the concentration of $P_{1,\psi_{1}(x)}$ and $P_{2,y_{2}}$,
$\langle f,\hat{S}f\rangle_{M_{1}}$ $=\int_{\mathsf{Y}_{1}}M_{1}(\mathrm{d}y_{1})\,f(y_{1})\int_{\mathsf{X}}P_{1,y_{1}}(\mathrm{d}x)\int_{\mathsf{X}}Q_{2,\psi_{2}(x)}(x,\mathrm{d}x^{\prime})\,f(\psi_{1}(x^{\prime}))$
$=\int_{\mathsf{X}}\Pi(\mathrm{d}x^{\prime\prime})\,f(\psi_{1}(x^{\prime\prime}))\int_{\mathsf{X}}P_{1,\psi_{1}(x^{\prime\prime})}(\mathrm{d}x)\int_{\mathsf{X}}Q_{2,\psi_{2}(x)}(x,\mathrm{d}x^{\prime})\,f(\psi_{1}(x^{\prime}))$
$=\int_{\mathsf{X}}\Pi(\mathrm{d}x)\,f(\psi_{1}(x^{\prime\prime}))\int_{\mathsf{X}}P_{1,\psi_{1}(x)}(\mathrm{d}x^{\prime\prime})\int_{\mathsf{X}}Q_{2,\psi_{2}(x)}(x,\mathrm{d}x^{\prime})\,f(\psi_{1}(x^{\prime}))$
$=\int_{\mathsf{X}}\Pi(\mathrm{d}x)\,f(\psi_{1}(x))\int_{\mathsf{X}}Q_{2,\psi_{2}(x)}(x,\mathrm{d}x^{\prime})\,f(\psi_{1}(x^{\prime}))$
$=\int_{\mathsf{Y}_{2}}M_{2}(\mathrm{d}y_{2})\int_{\mathsf{X}}P_{2,y_{2}}(\mathrm{d}x)\,f(\psi_{1}(x))\int_{\mathsf{X}}Q_{2,\psi_{2}(x)}(x,\mathrm{d}x^{\prime})\,f(\psi_{1}(x^{\prime}))$
$=\int_{\mathsf{Y}_{2}}M_{2}(\mathrm{d}y_{2})\int_{\mathsf{X}}P_{2,y_{2}}(\mathrm{d}x)\,f(\psi_{1}(x))\int_{\mathsf{X}}Q_{2,y_{2}}(x,\mathrm{d}x^{\prime})\,f(\psi_{1}(x^{\prime})).$
For $y_{2}\in\mathsf{Y}_{2}$, let $P_{2,y_{2}}(f\circ\psi_{1})=\int_{\mathsf{X}}P_{2,y_{2}}(\mathrm{d}x)\,f(\psi_{1}(x))$. Note that, for $M_{2}$-a.e. $y_{2}$, $P_{2,y_{2}}(f\circ\psi_{1})$ and $f\circ\psi_{1}-P_{2,y_{2}}(f\circ\psi_{1})$ can be viewed as functions in $L^{2}(P_{2,y_{2}})$ and $L_{0}^{2}(P_{2,y_{2}})$, respectively. Then by (9) in the proof of Theorem 9,
$\langle\hat{S}f,f\rangle_{M_{1}}=$ $\int_{\mathsf{Y}_{2}}M_{2}(\mathrm{d}y_{2})\,\langle f\circ\psi_{1}-P_{2,y_{2}}(f\circ\psi_{1}),Q_{2,y_{2}}[f\circ\psi_{1}-P_{2,y_{2}}(f\circ\psi_{1})]\rangle_{P_{2,y_{2}}}+$ (22)
$\int_{\mathsf{Y}_{2}}M_{2}(\mathrm{d}y_{2})\,\langle P_{2,y_{2}}(f\circ\psi_{1}),P_{2,y_{2}}(f\circ\psi_{1})\rangle_{P_{2,y_{2}}}.$
By (12) in the proof of Theorem 9, (1), and the concentration of $P_{1,y_{1}}(\cdot)$,
$\int_{\mathsf{Y}_{2}}M_{2}(\mathrm{d}y_{2})\,\langle P_{2,y_{2}}(f\circ\psi_{1}),P_{2,y_{2}}(f\circ\psi_{1})\rangle_{P_{2,y_{2}}}$ (23)
$=$ $\int_{\mathsf{X}}\Pi(\mathrm{d}x)\int_{\mathsf{X}}P_{2,\psi_{2}(x)}(\mathrm{d}x^{\prime})f(\psi_{1}(x^{\prime}))\,f(\psi_{1}(x))$
$=$ $\int_{\mathsf{Y}_{1}}M_{1}(\mathrm{d}y_{1})\int_{\mathsf{X}}P_{1,y_{1}}(\mathrm{d}x)\int_{\mathsf{X}}P_{2,\psi_{2}(x)}(\mathrm{d}x^{\prime})f(\psi_{1}(x^{\prime}))\,f(\psi_{1}(x))$
$=$ $\int_{\mathsf{Y}_{1}}M_{1}(\mathrm{d}y_{1})\int_{\mathsf{X}}P_{1,y_{1}}(\mathrm{d}x)\int_{\mathsf{X}}P_{2,\psi_{2}(x)}(\mathrm{d}x^{\prime})f(\psi_{1}(x^{\prime}))\,f(y_{1})$
$=$ $\langle f,Sf\rangle_{M_{1}}.$
Combining (22) and (23) gives
$\langle f,\hat{S}f\rangle_{M_{1}}=\int_{\mathsf{Y}_{2}}M_{2}(\mathrm{d}y_{2})\,\langle f\circ\psi_{1}-P_{2,y_{2}}(f\circ\psi_{1}),Q_{2,y_{2}}[f\circ\psi_{1}-P_{2,y_{2}}(f\circ\psi_{1})]\rangle_{P_{2,y_{2}}}+\langle f,Sf\rangle_{M_{1}}.$ (24)
By assumption, for $M_{2}$-a.e. $y_{2}\in\mathsf{Y}_{2}$,
$C_{L}\|f\circ\psi_{1}-P_{2,y_{2}}(f\circ\psi_{1})\|_{P_{2,y_{2}}}^{2}$ $\leq\langle f\circ\psi_{1}-P_{2,y_{2}}(f\circ\psi_{1}),Q_{2,y_{2}}[f\circ\psi_{1}-P_{2,y_{2}}(f\circ\psi_{1})]\rangle_{P_{2,y_{2}}}$
$\leq C_{U}\|f\circ\psi_{1}-P_{2,y_{2}}(f\circ\psi_{1})\|_{P_{2,y_{2}}}^{2}.$
Meanwhile, by (13) in the proof of Theorem 9 and then (23),
$\int_{\mathsf{Y}_{2}}\|f\circ\psi_{1}-P_{2,y_{2}}(f\circ\psi_{1})\|_{P_{2,y_{2}}}^{2}\,M_{2}(\mathrm{d}y_{2})$
$=$ $\|f\circ\psi_{1}\|_{\Pi}^{2}-\int_{\mathsf{X}}\Pi(\mathrm{d}x)\int_{\mathsf{X}}P_{2,\psi_{2}(x)}(\mathrm{d}x^{\prime})\,f(\psi_{1}(x^{\prime}))\,f(\psi_{1}(x))$
$=$ $\|f\|_{M_{1}}^{2}-\langle f,Sf\rangle_{M_{1}}.$
Thus,
$C_{L}(\|f\|_{M_{1}}^{2}-\langle f,Sf\rangle_{M_{1}})\leq$ $\int_{\mathsf{Y}_{2}}M_{2}(\mathrm{d}y_{2})\,\langle f\circ\psi_{1}-P_{2,y_{2}}(f\circ\psi_{1}),Q_{2,y_{2}}[f\circ\psi_{1}-P_{2,y_{2}}(f\circ\psi_{1})]\rangle_{P_{2,y_{2}}}$ (25)
$\leq$ $C_{U}(\|f\|_{M_{1}}^{2}-\langle f,Sf\rangle_{M_{1}}).$
Combining (24) and (25) yields (16). ∎
Appendix C Proof of Theorem 15
Before proceeding to the proof, we first give a useful representation of the operator $\hat{S}$. Define a linear operator $Q_{2}:L_{0}^{2}(\Pi)\to L_{0}^{2}(\Pi)$ as follows: for $f\in L_{0}^{2}(\Pi)$,
$Q_{2}f(x)=\int_{\mathsf{X}}Q_{2,\psi_{2}(x)}(x,\mathrm{d}x^{\prime})\,f(x^{\prime}),\quad x\in\mathsf{X}.$
[…]
$V_{1}f(x)=f(\psi_{1}(x)),\quad x\in\mathsf{X}.$
Define the linear transformation $V_{1}^{*}:L_{0}^{2}(\Pi)\to L_{0}^{2}(M_{1})$ as follows: for $f\in L_{0}^{2}(\Pi)$,
$V_{1}^{*}f(y_{1})=\int_{\mathsf{X}}f(x)\,P_{1,y_{1}}(\mathrm{d}x),\quad y_{1}\in\mathsf{Y}_{1}.$
Then $V_{1}^{*}$ is the adjoint of $V_{1}$ since, for $f\in L_{0}^{2}(\Pi)$ and $g\in L_{0}^{2}(M_{1})$,
$\langle f,V_{1}g\rangle_{\Pi}$ $=\int_{\mathsf{X}}f(x)\,g(\psi_{1}(x))\,\Pi(\mathrm{d}x)$
$=\int_{\mathsf{Y}_{1}}\int_{\mathsf{X}}f(x)\,g(\psi_{1}(x))\,P_{1,y_{1}}(\mathrm{d}x)\,M_{1}(\mathrm{d}y_{1})$
$=\int_{\mathsf{Y}_{1}}g(y_{1})\int_{\mathsf{X}}f(x)\,P_{1,y_{1}}(\mathrm{d}x)\,M_{1}(\mathrm{d}y_{1})$
$=\langle V_{1}^{*}f,g\rangle_{M_{1}}.$
Note that we have used (1) and the concentration of $P_{1,y_{1}}(\cdot)$ on $\psi^{-1}_{1}(y_{1})$.
Lemma 31.
Let $t$ be a positive integer. Note that, for $M_{i}$-a.e. $y_{i}\in\mathsf{Y}_{i}$, $Q_{i,y_{i}}^{t}(\cdot,\cdot)$ denotes the $t$-step Mtk of the chain associated with $Q_{i,y_{i}}(\cdot,\cdot)$. Then the operator associated with the Mtk $(x,A)\mapsto Q_{2,\psi_{2}(x)}^{t}(x,A)$ is the self-adjoint operator $Q_{2}^{t}$.
Proof.
Let $t$ be a positive integer. It suffices to show that, for $f\in L_{0}^{2}(\Pi)$,
$Q_{2}^{t}f(x)=\int_{\mathsf{X}}Q_{2,\psi_{2}(x)}^{t}(x,\mathrm{d}x^{\prime})\,f(x^{\prime}),\quad\Pi\text{-a.e. }x\in\mathsf{X}.$
To do this, use induction on $t$. The result holds when $t=1$ by definition. Assume that it holds for some positive integer $t$. Then, for $f\in L_{0}^{2}(\Pi)$ and $\Pi$-a.e. $x\in\mathsf{X}$,
$Q_{2}^{t+1}f(x)$ $=Q_{2}(Q_{2}^{t}f)(x)$
$=\int_{\mathsf{X}}\,Q_{2,\psi_{2}(x)}(x,\mathrm{d}x^{\prime})\int_{\mathsf{X}}Q_{2,\psi_{2}(x^{\prime})}^{t}(x^{\prime},\mathrm{d}x^{\prime\prime})\,f(x^{\prime\prime})$
$=\int_{\mathsf{X}}Q_{2,\psi_{2}(x)}(x,\mathrm{d}x^{\prime})\int_{\mathsf{X}}Q_{2,\psi_{2}(x)}^{t}(x^{\prime},\mathrm{d}x^{\prime\prime})\,f(x^{\prime\prime})$
$=\int_{\mathsf{X}}Q_{2,\psi_{2}(x)}^{t+1}(x,\mathrm{d}x^{\prime})f(x^{\prime}),$
where the third equality holds since, by Lemma 29, for $\Pi$-a.e. $x\in\mathsf{X}$, $Q_{2,\psi_{2}(x)}(x,\cdot)$ is concentrated on $\psi^{-1}_{2}(\psi_{2}(x))$. The desired result thus holds for all $t$. ∎
An immediate consequence of Lemma 31 is that the Mtk given by
$\hat{S}_{t}(y_{1},A)=\int_{\mathsf{X}}P_{1,y_{1}}(\mathrm{d}x)\,Q_{2,\psi_{2}(x)}^{t}(x,\psi_{1}^{-1}(A)),\quad y_{1}\in\mathsf{Y}_{1},\;A\in\mathcal{F}_{1},$ (26)
corresponds to the operator $V_{1}^{*}Q_{2}^{t}V_{1}$ for any positive integer $t$. In particular, $\hat{S}=\hat{S}_{1}=V_{1}^{*}Q_{2}V_{1}$. Another lemma that we need is the following.
Lemma 32.
Suppose that $Q_{2,y_{2}}$ is positive semi-definite for $M_{2}$-a.e. $y_{2}\in\mathsf{Y}_{2}$. Then $Q_{2}$ is positive semi-definite.
Proof.
It was already mentioned that $Q_{2}$ is self-adjoint. Let $f\in L_{0}^{2}(\Pi)$. By (1) and the concentration of $P_{2,y_{2}}(\cdot)$,
$\langle f,Q_{2}f\rangle_{\Pi}$ $=\int_{\mathsf{X}}\Pi(\mathrm{d}x)f(x)\int_{\mathsf{X}}Q_{2,\psi_{2}(x)}(x,\mathrm{d}x^{\prime})\,f(x^{\prime})$
$=\int_{\mathsf{Y}_{2}}M_{2}(\mathrm{d}y_{2})\int_{\mathsf{X}}P_{2,y_{2}}(\mathrm{d}x)f(x)\int_{\mathsf{X}}Q_{2,y_{2}}(x,\mathrm{d}x^{\prime})\,f(x^{\prime}).$
Let $P_{2,y_{2}}f=\int_{\mathsf{X}}P_{2,y_{2}}(\mathrm{d}x)\,f(x)$. For $M_{2}$-a.e. $y_{2}\in\mathsf{Y}_{2}$, $P_{2,y_{2}}f$ can be regarded as a constant function in $L^{2}(P_{2,y_{2}})$, and $f-P_{2,y_{2}}f$ can be regarded as a function in $L_{0}^{2}(P_{2,y_{2}})$. Then
$\int_{\mathsf{X}}P_{2,y_{2}}(\mathrm{d}x)f(x)\int_{\mathsf{X}}Q_{2,y_{2}}(x,\mathrm{d}x^{\prime})\,f(x^{\prime})$
$=$ $\langle f-P_{2,y_{2}}f,Q_{2,y_{2}}(f-P_{2,y_{2}}f)\rangle_{P_{2,y_{2}}}+\langle P_{2,y_{2}}f,Q_{2,y_{2}}(f-P_{2,y_{2}}f)\rangle_{P_{2,y_{2}}}+$
$\langle f-P_{2,y_{2}}f,P_{2,y_{2}}f\rangle_{P_{2,y_{2}}}+(P_{2,y_{2}}f)^{2}$
$=$ $\langle f-P_{2,y_{2}}f,Q_{2,y_{2}}(f-P_{2,y_{2}}f)\rangle_{P_{2,y_{2}}}+(P_{2,y_{2}}f)^{2}\geq 0.$
It follows that $\langle f,Q_{2}f\rangle_{\Pi}\geq 0$, and thus $Q_{2}$ is positive semi-definite. ∎
Recall the statement of Theorem 15. Theorem 15. Let $\gamma:\mathsf{Y}_{2}\to[0,1]$ be a measurable function such that $\|Q_{2,y_{2}}\|_{P_{2,y_{2}}}\leq\gamma(y_{2})$ for $M_{2}$-a.e. $y_{2}\in\mathsf{Y}_{2}$. Let $t$ be some positive integer, and let $\alpha_{t}\in[0,1]$ be such that
$\int_{\mathsf{X}}\gamma(\psi_{2}(x))^{t}\,P_{1,y_{1}}(\mathrm{d}x)\leq\alpha_{t}$ (27)
for $M_{1}$-a.e. $y_{1}\in\mathsf{Y}_{1}$. If either (i) $t$ is even, or (ii) $Q_{2,y_{2}}$ is positive semi-definite for $M_{2}$-a.e. $y_{2}\in\mathsf{Y}_{2}$, then, for $f\in L_{0}^{2}(M_{2})\setminus\{0\}$,
$0\leq\frac{\langle f,\hat{S}f\rangle_{M_{1}}^{t}}{\|f\|_{M_{1}}^{2t}}\leq\frac{\langle f,Sf\rangle_{M_{1}}}{\|f\|_{M_{1}}^{2}}+\alpha_{t},$ (18)
i.e.,
$1-\frac{\langle f,Sf\rangle_{M_{1}}}{\|f\|_{M_{1}}^{2}}-\alpha_{t}\leq 1\frac{\langle f,\hat{S}f\rangle_{M_{1}}^{t}}{\|f\|_{M_{1}}^{2t}}\leq 1$
Proof.
Let $t$ be a positive integer for which (27) holds, and let $f\in L_{0}^{2}(M_{1})$ be nonzero. Let $\hat{S}_{t}$ be defined as in (26). Replacing $Q_{2,y_{2}}$ with $Q_{2,y_{2}}^{t}$ in (24) in the proof of Theorem 12 yields
$\langle f,\hat{S}_{t}f\rangle_{M_{1}}=\int_{\mathsf{Y}_{2}}M_{2}(\mathrm{d}y_{2})\,\langle f\circ\psi_{1}-P_{2,y_{2}}(f\circ\psi_{1}),Q_{2,y_{2}}^{t}[f\circ\psi_{1}-P_{2,y_{2}}(f\circ\psi_{1})]\rangle_{P_{2,y_{2}}}+\langle f,Sf\rangle_{M_{1}}.$
By reversibility, $\|Q_{2,y_{2}}^{t}\|_{P_{2,y_{2}}}=\|Q_{2,y_{2}}\|_{P_{2,y_{2}}}^{t}$ holds almost everywhere, so
$\langle f,\hat{S}_{t}f\rangle_{M_{1}}\leq\int_{\mathsf{Y}_{2}}M_{2}(\mathrm{d}y_{2})\,\gamma(y_{2})^{t}\|f\circ\psi_{1}-P_{2,y_{2}}(f\circ\psi_{1})\|_{P_{2,y_{2}}}^{2}+\langle f,Sf\rangle_{M_{1}}.$
The variance of a random variable is no greater than the second moment, so
$\langle f,\hat{S}_{t}f\rangle_{M_{1}}$ $\leq\int_{\mathsf{Y}_{2}}M_{2}(\mathrm{d}y_{2})\,\gamma(y_{2})^{t}\|f\circ\psi_{1}\|_{P_{2,y_{2}}}^{2}+\langle f,Sf\rangle_{M_{1}}$
$=\int_{\mathsf{Y}_{2}}M_{2}(\mathrm{d}y_{2})\,\gamma(y_{2})^{t}\int_{\mathsf{X}}f(\psi_{1}(x))^{2}\,P_{2,y_{2}}(\mathrm{d}x)+\langle f,Sf\rangle_{M_{1}}$
$=\int_{\mathsf{Y}_{2}}\int_{\mathsf{X}}M_{2}(\mathrm{d}y_{2})\,P_{2,y_{2}}(\mathrm{d}x)\,\gamma(\psi_{2}(x))^{t}f(\psi_{1}(x))^{2}+\langle f,Sf\rangle_{M_{1}}$
$=\int_{\mathsf{X}}\Pi(\mathrm{d}x)\,\gamma(\psi_{2}(x))^{t}f(\psi_{1}(x))^{2}+\langle f,Sf\rangle_{M_{1}}$
$=\int_{\mathsf{Y}_{1}}\int_{\mathsf{X}}M_{1}(\mathrm{d}y_{1})\,P_{1,y_{1}}(\mathrm{d}x)\,\gamma(\psi_{2}(x))^{t}f(\psi_{1}(x))^{2}+\langle f,Sf\rangle_{M_{1}}$
$=\int_{\mathsf{Y}_{1}}\int_{\mathsf{X}}M_{1}(\mathrm{d}y_{1})\,P_{1,y_{1}}(\mathrm{d}x)\,\gamma(\psi_{2}(x))^{t}f(y_{1})^{2}+\langle f,Sf\rangle_{M_{1}}$
$\leq\alpha_{t}\|f\|_{M_{1}}^{2}+\langle f,Sf\rangle_{M_{1}},$
where we have used (1) and the concentration of $P_{2,y_{2}}(\cdot)$ and $P_{1,y_{1}}(\cdot)$. Assume that $t$ is even, or that $Q_{2,y_{2}}$ is positive semi-definite for $M_{2}$-a.e. $y_{2}\in\mathsf{Y}_{2}$. Then, by Lemma 32, $t$ is even, or $Q_{2}$ is positive semi-definite. Note also that $\|V_{1}f\|_{\Pi}^{2}=\|f\|_{M_{1}}^{2}$. By Lemmas 30 and 31,
$0\leq\left(\frac{\langle f,\hat{S}f\rangle_{M_{1}}}{\|f\|_{M_{1}}^{2}}\right)^{t}=\left(\frac{\langle V_{1}f,Q_{2}V_{1}f\rangle_{\Pi}}{\|V_{1}f\|_{\Pi}^{2}}\right)^{t}\leq\frac{\langle V_{1}f,Q_{2}^{t}V_{1}f\rangle_{\Pi}}{\|V_{1}f\|_{\Pi}}=\frac{\langle f,\hat{S}_{t}f\rangle_{M_{1}}}{\|f\|_{M_{1}}^{2}}.$
Combining the previous two displays yields (18). ∎
Appendix D Proof of Proposition 17
Recall the statement of Proposition 17. Proposition 17. In the context of Examples 1 and 3, for $i=1,\dots,n$ and $M_{i}$-a.e. $y_{i}\in\mathsf{Y}_{i}$, there exists an isometric isomorphism $U_{i,y_{i}}:L_{0}^{2}(P_{i,y_{i}})\to L_{0}^{2}(\Pi_{i,y_{i}})$ such that $Q_{i,y_{i}}=U_{i,y_{i}}^{-1}R_{i,y_{i}}U_{i,y_{i}}$. As a consequence, $Q_{i,y_{i}}$ is positive semi-definite if and only if $R_{i,y_{i}}$ is, and $\|Q_{i,y_{i}}\|_{P_{i,y_{i}}}=\|R_{i,y_{i}}\|_{\Pi_{i,y_{i}}}$.
Proof.
Fix $i=1,\dots,n$ and $y_{i}\in\mathsf{Y}_{i}$ such that $\Pi_{i,y_{i}}$ is a probability measure (in case the conditional probability is well-defined only almost surely). For $f\in L_{0}^{2}(P_{i,y_{i}})$, let $U_{i,y_{i}}f$ be the function on $\mathsf{X}_{i}$ such that $U_{i,y_{i}}f(z)=f(x_{1},\dots,x_{n})$ whenever $x_{i}=z\in\mathsf{X}_{i}$ and $x_{-i}=y_{i}$. Then $\Pi_{i,y_{i}}(U_{i,y_{i}}f)=P_{i,y_{i}}f$, and
$\int_{\mathsf{X}_{i}}[U_{i,y_{i}}f(x_{i})]^{2}\,\Pi_{i,y_{i}}(\mathrm{d}x_{i})=\int_{\mathsf{X}}f(x)^{2}\,P_{i,y_{i}}(\mathrm{d}x).$
This implies that $U_{i,y_{i}}$ is a transformation from $L_{0}^{2}(P_{i,y_{i}})$ to $L_{0}^{2}(\Pi_{i,y_{i}})$, and $\|U_{i,y_{i}}f\|_{\Pi_{i,y_{i}}}=\|f\|_{P_{i,y_{i}}}$ for $f\in L_{0}^{2}(P_{i,y_{i}})$. The transformation is clearly linear. To prove that it is an isometric isomorphism, we only need to show that it is bijective. It is injective since if $U_{i,y_{i}}f=0$ for some $f\in L_{0}^{2}(P_{i,y_{i}})$ then $\|f\|_{P_{i,y_{i}}}=\|U_{i,y_{i}}f\|_{\Pi_{i,y_{i}}}=0$. To see that it surjective, note that, for $g\in L_{0}^{2}(\Pi_{i,y_{i}})$, one can identify a function $f\in L_{0}^{2}(P_{i,y_{i}})$ such that $f(x_{1},\dots,x_{n})=g(z)$ whenever $x_{i}=z\in\mathsf{X}_{i}$, which implies that $U_{i,y_{i}}f=g$. It is evident that, for $f\in L_{0}^{2}(P_{i,y_{i}})$ and $z\in\mathsf{X}_{i}$,
$U_{i,y_{i}}Q_{i,y_{i}}f(z)=\int_{\mathsf{X}}R_{i,y_{i}}(z,\mathrm{d}x_{i})\,\delta_{y_{i}}(\mathrm{d}x_{-i})\,f(x_{1},\dots,x_{n})=R_{i,y_{i}}U_{i,y_{i}}f(z).$
Thus, $Q_{i,y_{i}}=U_{i,y_{i}}^{-1}R_{i,y_{i}}U_{i,y_{i}}$. The final claim of the proposition holds since, for $f\in L_{0}^{2}(P_{i,y_{i}})\setminus\{0\}$,
$\frac{\langle f,Q_{i,y_{i}}f\rangle_{P_{i,y_{i}}}}{\|f\|_{P_{i,y_{i}}}^{2}}=\frac{\langle U_{i,y_{i}}f,R_{i,y_{i}}U_{i,y_{i}}f\rangle_{\Pi_{i,y_{i}}}}{\|U_{i,y_{i}}f\|_{\Pi_{i,y_{i}}}^{2}}.$
Here, we used the fact that the inverse of an isometric isomorphism is its adjoint. ∎
Appendix E Proof of Lemma 25
Recall the statement of the lemma. Lemma 25. Suppose that (A) holds. Let $\Lambda\subset[k]$ and $\Lambda^{\prime}\subset\Lambda$ be such that |Λ|=normal-$|\Lambda|=\ell$ and |Λ|=superscriptnormal-$|\Lambda^{\prime}|=\ell^{\prime}$ (where $\ell>\ell^{\prime}$). Then $P_{\Lambda,\psi_{\Lambda}(x);\,\Lambda^{\prime},\psi_{\Lambda^{\prime}}(x)}(\cdot)=P_{\Lambda^{\prime},\psi_{\Lambda^{\prime}}(x)}(\cdot)$ for $\Pi$-a.e. $x\in\mathsf{X}$.
Proof.
By the definition of disintegration, for $y_{\Lambda}\in\mathsf{Y}_{\Lambda}$ and $P_{\Lambda,y_{\Lambda}}$-a.e. $x\in\mathsf{X}$, $P_{\Lambda,y_{\Lambda};\,\Lambda^{\prime},\psi_{\Lambda^{\prime}}(x)}$ is concentrated on $\psi^{-1}_{\Lambda^{\prime}}(\psi_{\Lambda^{\prime}}(x))$. It then follows from (1) and the concentration of $P_{\Lambda,y_{\Lambda}}$ that
$\int_{\mathsf{X}}\Pi(\mathrm{d}x)\,P_{\Lambda,\psi_{\Lambda}(x);\,\Lambda^{\prime},\psi_{\Lambda^{\prime}}(x)}(\mathsf{X}\setminus\psi^{-1}_{\Lambda^{\prime}}(\psi_{\Lambda^{\prime}}(x)))$
$=$ $\int_{\mathsf{Y}_{\Lambda}}M_{\Lambda}(\mathrm{d}y_{\Lambda})\,\int_{\mathsf{X}}P_{\Lambda,y_{\Lambda}}(\mathrm{d}x)\,P_{\Lambda,\psi_{\Lambda}(x);\,\Lambda^{\prime},\psi_{\Lambda^{\prime}}(x)}(\mathsf{X}\setminus\psi^{-1}_{\Lambda^{\prime}}(\psi_{\Lambda^{\prime}}(x)))$
$=$ $\int_{\mathsf{Y}_{\Lambda}}M_{\Lambda}(\mathrm{d}y_{\Lambda})\,\int_{\mathsf{X}}P_{\Lambda,y_{\Lambda}}(\mathrm{d}x)\,P_{\Lambda,y_{\Lambda};\,\Lambda^{\prime},\psi_{\Lambda^{\prime}}(x)}(\mathsf{X}\setminus\psi^{-1}_{\Lambda^{\prime}}(\psi_{\Lambda^{\prime}}(x)))$
$=$ $0$
The integrals are well-defined because the function
[…]
where $A_{\Lambda}=\{(x,x^{\prime})\in\mathsf{X}^{2}:\,\psi_{\Lambda}(x^{\prime})\neq\psi_{\Lambda}(x)\}$, is measurable by (H) and (A). In conclusion, for $\Pi$-a.e. $x\in\mathsf{X}$, $P_{\Lambda,\psi_{\Lambda}(x);\,\Lambda^{\prime},\psi_{\Lambda^{\prime}}(x)}$ is concentrated on $\psi^{-1}_{\Lambda}(\psi_{\Lambda^{\prime}}(x))$. On the other hand, for $A\in\mathcal{B}$,
$\int_{\mathsf{X}}\Pi(\mathrm{d}x)\,P_{\Lambda,\psi_{\Lambda}(x);\,\Lambda^{\prime},\psi_{\Lambda^{\prime}}(x)}(A)$ $=\int_{Y_{\Lambda}}M_{\Lambda}(\mathrm{d}y_{\Lambda})\int_{\mathsf{X}}P_{\Lambda,y_{\Lambda}}(\mathrm{d}x)\,P_{\Lambda,\psi_{\Lambda}(x);\,\Lambda^{\prime},\psi_{\Lambda^{\prime}}(x)}(A)$
$=\int_{Y_{\Lambda}}M_{\Lambda}(\mathrm{d}y_{\Lambda})\int_{\mathsf{X}}P_{\Lambda,y_{\Lambda}}(\mathrm{d}x)\,P_{\Lambda,y_{\Lambda};\,\Lambda^{\prime},\psi_{\Lambda^{\prime}}(x)}(A)$
$=\int_{Y_{\Lambda}}M_{\Lambda}(\mathrm{d}y_{\Lambda})\,P_{\Lambda,y_{\Lambda}}(A)$
$=\Pi(A).$
The third equality follows from (1) with $\Pi$ replaced by $P_{\Lambda,y_{\Lambda}}$ and $P_{i,y_{i}}$ replaced by $P_{\Lambda,y_{\Lambda};\,\Lambda^{\prime},z_{\Lambda^{\prime}}}$. For $z_{\Lambda^{\prime}}\in\mathsf{Y}_{\Lambda^{\prime}}$, let $\tilde{P}_{\Lambda^{\prime},z_{\Lambda^{\prime}}}$ be the measure $P_{\Lambda,\psi_{\Lambda}(x);\,\Lambda^{\prime},\psi_{\Lambda^{\prime}}(x)}$ if $\psi_{\Lambda^{\prime}}(x)=z_{\Lambda^{\prime}}$. Note that $\tilde{P}_{\Lambda^{\prime},z_{\Lambda^{\prime}}}$ is uniquely defined because $\psi_{\Lambda}(x)=(x_{i})_{i\not\in\Lambda}$ depends on $x$ only through $\psi_{\Lambda^{\prime}}(x)=(x_{i})_{i\not\in\Lambda^{\prime}}$. By previous analyses, $(\tilde{P}_{\Lambda^{\prime},z_{\Lambda^{\prime}}})_{z_{\Lambda^{\prime}}\in\mathsf{Y}_{\Lambda^{\prime}}}$ is a disintegration of $\Pi$ with respect to $\psi_{\Lambda^{\prime}}$. Because $\mathsf{X}$ and $\mathsf{Y}_{\Lambda}$ are both Polish, the disintegration is unique up to $M_{\Lambda^{\prime}}$ almost sure equivalence (see, e.g., Chang and Pollard, 1997, Theorem 1). In particular, For $M_{\Lambda}$-a.e. $z_{\Lambda^{\prime}}\in\mathsf{Y}_{\Lambda^{\prime}}$, $\tilde{P}_{\Lambda^{\prime},z_{\Lambda^{\prime}}}=P_{\Lambda^{\prime},z_{\Lambda^{\prime}}}$. This proves the desired result. ∎